### v3

# Towards Bidirectional Human-AI Alignment: A Systematic Review for Clarifications, Framework, and Future Directions

Abstract.
Recent advancements in general-purpose AI have highlighted the importance of guiding AI systems towards the intended goals, ethical principles, and values of individuals and groups, a concept broadly recognized as alignment. However, the lack of clarified definitions and scopes of human-AI alignment poses a significant obstacle, hampering collaborative efforts across research domains to achieve this alignment. In particular, ML- and philosophy-oriented alignment research often views AI alignment as a static, unidirectional process (i.e., aiming to ensure that AI systems’ objectives match humans) rather than an ongoing, mutual alignment problem (Wikipedia, 2024). This perspective largely neglects the long-term interaction and dynamic changes of alignment. To understand these gaps, we introduce a systematic review of over 400 papers published between 2019 and January 2024, spanning multiple domains such as Human-Computer Interaction (HCI), Natural Language Processing (NLP), Machine Learning (ML). We characterize, define and scope human-AI alignment. From this, we present a conceptual framework of “Bidirectional Human-AI Alignment” to organize the literature from a human-centered perspective. This framework encompasses both 1) conventional studies of aligning AI to humans that ensures AI produces the intended outcomes determined by humans, and 2) a proposed concept of aligning humans to AI, which aims to help individuals and society adjust to AI advancements both cognitively and behaviorally. Additionally, we articulate the key findings derived from literature analysis, including literature gaps and trends, human values, and interaction techniques. To pave the way for future studies, we envision three key challenges and give recommendations for future research.
Human-AI Alignment, Human-AI Interaction
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Manuscript; submitted to ACM; 2024ccs: Human-centered computing HCI theory, concepts and models
1. Introduction
Artificial Intelligence (AI) has advanced significantly, especially with the advent of general-purpose generative AI, demonstrating unprecedented capabilities in solving a wide range of complicated and challenging problems, such as reasoning, generation, language understanding, and more (Ouyang et al., 2022). However, as AI becomes increasingly powerful and integrated into daily life, society is confronted with an array of risks to both individuals and to broader society (Morris et al., 2024). For example, text-to-image generative models were found to amplify stereotypes about race and gender (blo, 2023), and biased algorithms in hiring processes were found to perpetuate discrimination (Park et al., 2021a). These risks illuminate foundational questions around how and which values are embedded in AI models that drive decisions within real-world contexts.
[…]
Figure 1. The overview of the Bidirectional Human-AI Alignment framework. Our framework encompasses both Refer to caption conventional studies of “Align AI to Humans” in AI development that ensures AI produces the intended outcomes determined by humans, and  Refer to caption a novel concept of “Align Humans to AI” , which aims to help humans and society to better understand, critique, collaborate, and adapt to transformative AI advancements. Note that “Humans” refers to both AI users and those who do not interact with AI but may be impacted by AI systems. We further identify four key research questions (i.e., RQ1-RQ4) to facilitate this holistic loop of “bidirectional human-AI alignment”, and organize the literature that can potentially address RQ1-RQ4 in Section 4.
This paper conducts a systematic literature review of over 400 human-AI alignment papers across multiple disciplines, based on PRISMA guidelines (Page et al., 2021; Stefanidi et al., 2023). The literature review corpus includes papers drawn from the the Human-Computer Interaction (HCI), Natural Language Processing (NLP) and Machine Learning (ML) domains that were published between the advent of general-purpose generative AI to present, i.e., primarily between January 2019 and January 2024. We take an interdisciplinary approach to human-AI alignment (Gilbert, 2024), drawing from theories and empirical studies across the aforementioned domains. To round out our approach, we composed our research team of members from the domains of HCI, NLP, ML, Data Science, Computational Social Science, and Cognitive Science.
    Definition. Bidirectional Human-AI Alignment is a comprehensive framework that encompasses two interconnected alignment processes: ‘Aligning AI to Humans’ and ‘Aligning Humans to AI’. The former focuses on integrating human specifications to train, steer, and customize AI, while the latter investigates human cognitive and behavioral adaptations to AI, which supports humans in understanding, critiquing, collaborating with, and adapting to AI advancements.
Paper Organization and Key Findings. We provide clarified definitions and scopes related to human-AI alignment, including “what is the goal of alignment?”, “with whom to align?”, and “what values should be aligned with?”(Section 3). We then present a conceptual framework of Bidirectional Human-AI Alignment from a long-term and dynamic perspective, encompassing both “Align AI to Humans” and “Align Humans to AI” (see Figure 1). Furthermore, we identify four key research questions (RQs) in the Bidirectional Human-AI Alignment framework and provide a structured way to organize the existing research literature to address these questions in Section 4. The resulting structured topologies in Figure 2 and 3 aim to provide a shared vocabulary that can help streamline communication and collaboration between alignment researchers in different disciplines. Furthermore, through iterative paper coding and literature analysis, we derive insights including meta analysis on gaps and trends in existing alignment research, findings of human values, and the interaction techniques for human-AI alignment (Section 5). To pave the way for future studies, we further envision three future directions in Figure 6 and Section 6. Our study highlights several key findings. For instance, many essential human values outlined in Table 2, such as loyalty and environmental protection, are crucial for alignment but are often overlooked in current research. Additionally, many topics within the bidirectional alignment framework, including leveraging implicit human feedback to align AI models and auditing AI for various social values, are important but underexplored.
Operationalizing the Framework. To help AI practitioners (e.g., developers, researchers, user experience designers) who are interested in this field to leverage the bidirectional approaches, we highlight some systematic resources about human-AI alignment provided both within this survey and on companion Github repository:
  • A Github repository of “Bidirectional Human-AI Alignment” maintaining latest reading lists and relevant updates.
  • Fig. 1: An overview of the bidirectional human-AI alignment framework with formal definition.
  • Table 1: Typical alignment goals, their definitions, and limitations.
  • Table 2: A comprehensive list of basic human values, their relationships, and related alignment papers.
  • Fig. 2 and 3: Fine-grained topologies and papers of “Aligning AI to Humans” & “Aligning Humans to AI” research.
  • Fig. 4: An illustration of trends and gaps of alignment research dimension distribution.
  • Fig. 5: A summary of common interactive techniques to specify human values and exemplary papers.
2. Challenges in Achieving Alignment
[…]
Challenge 1: Outer and Inner Alignment. In the context of “intelligent agents,” until now, AI alignment research has aimed to ensure that any AI systems that would be set free to make decisions on our behalf would act appropriately and reduce unintended consequences (Wooldridge, 1999; Stuurman and Wijnands, 2001; Russell and Norvig, 2016). At the near-term stage, aligning AI involves two main challenges: carefully specifying the purpose of the system (outer alignment i.e., providing well-specified rewards (Ngo et al., 2024)) and ensuring that the system adopts the specification robustly (inner alignment, i.e., ensuring that every action given an agent in a particular state learns desirable internally-represented goals (Ngo et al., 2024)). Significant efforts have been made, for inner alignment, to align AI systems to follow alignment goals of an individual or a group (e.g., instructions, preferences, values, and/or ethical principles) (Ouyang et al., 2022) and to evaluate the performance of alignment (Santurkar et al., 2023). However, for outer alignment, AI designers are still facing difficulties in specifying the full range of desired and undesired alignment goals of humans.
Challenge 2: Specification Gaming. To learn human alignment goals, AI designers typically provide an objective function, instructions, reward function, or feedback to the system, which is often unable to completely specify all important values and constraints that a human intended (Hemphill, 2020). Hence, AI designers resort to easy-to-specify proxy goals such as maximizing the approval of human overseers (Wikipedia, 2024), which results in “specification gaming” (Krakovna et al., 2020) or “reward hacking” (Pan et al., 2022) issues (i.e., AI systems can find loopholes that help them accomplish the specific objective efficiently but in unintended, possibly harmful ways). Additionally, the black-box nature of neural networks brings additional ethical and safety concerns for alignment because humans don’t know about the inner state and the actions AI leveraged to achieve the output. Consequently, AI systems might make “correct” decisions with “incorrect” reasons, which are difficult to discern. Society is already facing these issues, such as data privacy (Tucker et al., 2018), algorithmic bias (Hajian et al., 2016), self-driving car accidents (Bonnefon et al., 2016), and more. As a result, these considerations necessitate considering human-AI interaction in AI alignment for specification and evaluation, ranging from addressing problems around who uses an AI system, with what goals to specify, and if the AI system perform its intended function from the user’s perspective (Weidinger et al., 2023).
Challenge 3: Scalable Oversight. From a long-term perspective, when advanced AI systems become more complex and capable (e.g., AGI (Morris et al., 2024)), it becomes increasingly difficult to align them to human values through human feedback. Evaluating complex AI behaviors applied to increasingly challenging tasks can be slow or infeasible for humans to ensure all sub-steps are aligned with their values (Terry et al., 2023). Therefore, researchers have begun to investigate how to reduce the time and effort for human supervision, and how to assist human supervisors, referred to as Scalable Oversight (Amodei et al., 2016).
Challenge 4: Dynamic Nature. As AI systems become increasingly powerful, the alignment solutions must also adapt dynamically since human values and preferences change as well. As Dautenhahn et al. (2000) posit, AI systems may be neither humane nor desirable if we do not ask questions about the long-term cognitive and social effects of social agent systems (e.g., how will agent technology affect human cognition). All these considerations call for a long-term and dynamic perspective to address human-AI alignment as an ongoing, mutual process with the collective efforts of cross-domain expertise.
Challenge 5: Existential Risk. Further, some AI researchers claim that (Bengio et al., 2024) advanced AI systems will begin to seek power over their environment (e.g., humans) once deployed in real-world settings, as such behavior may not be noticed during training. For example, some language models seek power in text-based social environments by gaining money, resources, or social influence (Pan et al., 2023). Consequently, some hypothesize that future AI, if not properly aligned with human values, could pose an existential risk to humans (Dafoe and Russell, 2016).
Table 1. The Goals of Alignment. We present the six prevailing alignment goals, associating with their Definitions (middle column), Limitations and Risks (right column). We consider Human Values as the main goal of alignment in this work referring to an extensive analysis and arguments in existing studies (Gabriel, 2020; Russell, 2014)
3. Fundamental Definitions and Clarifications
Addressing alignment challenges is a complex and multifaceted process. To develop a comprehensive perspective of the ongoing and mutual process of human-AI alignment, this section begins by introducing core definitions and key components of alignment. Subsequently, we outline our methodology for conducting a systematic literature review and the rationale behind our proposed framework of bidirectional human-AI alignment.
3.1. What is the goal of alignment?
[…]
Table 2. The relations and a fine-grained taxonomy of 69 exemplary human values. We consider 5 high-order value types encompassing 12 motivational value types, indicated by their sources (e.g., individuals, society and interaction). The exemplary values with red dot ([Uncaptioned image]) indicates there are no work in our surveyed papers examining the specific values.
3.2. With whom to align?
Pertinent stakeholders within the AI landscape can be the potential objects for AI to align with, including lay people (e.g., end users of AI systems) (Qi et al., 2023), AI practitioners (e.g., developers, researchers) (Boggust et al., 2022; Anwar et al., 2024; Ngo et al., 2024), organizational entities (e.g., technology firms, professional communities), national/international bodies (e.g., governments, legislative bodies) (Deshpande and Sharp, 2022) and others. Many alignment research papers have focused on general humans without specifying particular groups (Yao et al., 2023b; Liu et al., 2022; Guo et al., 2024). Nevertheless, different groups hold different, sometimes even contrasting, values (Gabriel, 2020). As a consequence, rather than identifying a true moral theory as a one-size-fits-all value, prior alignment research argues to select the appropriate principles for compatible human groups (Sorensen et al., 2024). To this end, pluralistic value alignment, grounded on social choice theory (Arrow, 2012), proposes combining individual views fairly in developing alignment principles as a potential solution (Sorensen et al., 2024). In this work, we also consider values from this pluralistic perspective, where AI should be aligned with pluralistic human individuals and societal groups who would ultimately be impacted by AI.
3.3. What are the values to be aligned with?
[…]
A Comprehensive Taxonomy of Human Values. Nevertheless, this conventional theory was developed without the context of human-AI interaction, which might overlook values that need to be considered for human-AI alignment. Therefore, we used a bottom-up approach to extract all values studied in our collected alignment literature (elaborated in Section A.1), mapped them onto the Schwartz Theory of Basic Values, and supplemented the theory with AI-related structure and content. As a result, we identified the structural relationships among human values (see Figure LABEL:fig:value_structure) and mapped existing literature to a fine-grained taxonomy (see Table 2). As shown in Figure LABEL:fig:value_structure, we supplemented the traditional theory’s four high-order value types (i.e., “Self-Enhancement”, “Openness to Change”, “Conservation”, “Self-Transcendence”) with a novel high-order value type, named “Desired Values for AI Tools” that encompasses two motivational value types (i.e., “Usability” and “Human-Likeness”). We further organize the relationship among these value types along two dimensions (Schwartz, 2012): different resources (i.e., individuals, society and interaction) and different self-intentions (i.e., self-protection against threat and self-expansion and growth). Furthermore, we elaborate the definitions of the 12 motivational value types and their exemplary values by mapping them to relevant human-AI alignment papers from our corpus in Table 2. During the process of mapping, we found: 1) value terms in empirical papers were often named differently (e.g., capability and competence), or check their opposites (e.g., fairness and bias); 2) there are many values not studied in our corpus, i.e., indicated as ([Uncaptioned image]) in the Figure.
    Takeaways & Implications for Alignment Clarifications 1. We propose the “pluralistic human values” as the alignment goal, wherein AI should be aligned with the diverse perspectives of individuals and societal groups who may be directly or indirectly impacted by AI. 2. Many exemplary human values outlined in Table 2, such as loyalty and environmental protection, are important for alignment but are often overlooked or underexplored in current research. 3. The human value system should account for complex application contexts and relationships, such as priorities and conflicts, rather than focusing on an exclusive subset of values.
3.4. Scopes and Key Components in Alignment
This section clarifies the key components we considered in human-AI alignment. Additionally, we define the scopes of each key component as commonly adopted in current research.
  • Humans. Our primary focus is on human individuals, groups, or organizations that will ultimately develop, use, and potentially impact or be influenced by AI systems, as these are the entities with which AI systems should align. We emphasize the importance of considering the pluralistic values of diverse users rather than treating users as a monolithic group.
  • Artificial Intelligence (AI). We focus on AI systems including both domain-specific AI systems that address specific tasks (e.g., reasoning, dialogue) and general-purpose AI models that aim to complete any tasks with performance comparable to a human’s. These AI systems include generative, classification, and regression models, among others. Particularly, we primarily focus on language models as the representative AI models for alignment research, and discuss how the insights from this study can be generalized to other modalities at the section end.
  • Alignment. Our review encompassed all the alignment goals outlined in Section 3.1. Since many studies emphasized the importance of value alignment (Gabriel, 2020; Russell, 2014), we particularly summarized a clarified taxonomy of alignment values and identified the value in each paper (if applicable). Additionally, we focus on analyzing the AI models’ output and generation, but not the neural network’s intermediate representations (McGrath et al., 2023; Park et al., 2023; Xie et al., 2024), for alignment research.
3.5. Systematic Literature Review and Framework Rationale
Achieving the alignment between AI and human values necessitates collaboration among interdisciplinary AI practitioners (e.g., developers and researchers). To comprehensively analyze the research literature pertaining to this ongoing, mutual process of human-AI alignment, we conducted a systematic literature review adhering to the the PRISMA guideline (Page et al., 2021; Stefanidi et al., 2023). Our iterative selection process yielded 411 papers from an initial pool of 34,213 publications, spanning the period from January 2019 to January 2024. These papers were sourced from leading venues across multiple AI-related domains, including Natural Language Processing, Human-Computer Interaction, Machine Learning, and more, using well-defined keywords and criteria. A group of 13 interdisciplinary researchers coded the papers. Through iterative discussions, we developed the paper codes and proposed framework (Section 3). Detailed steps and processes are elucidated in Appendix A.
3.6. Rationale Behind the Framework
Our systematic review revealed that current research largely overlooks the nature of long-term dynamic changes and the importance of interaction in human-AI alignment. First, the manifestation of certain AI risks, harms, and the practical quality of AI services primarily emerge through interaction between humans and AI systems (Weidinger et al., 2023). For instance, Shen and Huang (2020) discovered that AI explanations, deemed effective by AI developers and researchers, were found not useful for humans in real-world scenarios of interactive AI error detection. Additionally, the use of AI influences human thinking and behavior, which in turn alters their values when interacting with AI. For instance, Li et al. (2024b) found that incorporating AI assistance in writing workflows changed humans’ writing perceptions, behavior, and performance, creating new requirements for developing the next generation of AI writing assistants (Lee et al., 2024).
In light of these findings, we posit that AI practitioners must expand their focus beyond merely “Aligning AI to Humans” – a paradigm centered on integrating human specifications to train, steer, and customize AI. Equal emphasis must be placed on “Aligning Humans to AI”, a complementary paradigm investigating human cognitive and behavioral adaptation to AI. This latter aspect is crucial in supporting individuals and society at large in understanding, critiquing, collaborating with, and adapting to transformative AI advancements. To encapsulate these comprehensive perspectives, we propose the “Bidirectional Human-AI Alignment” framework, and provide its definition below Figure 1. This holistic framework acknowledges the reciprocal nature of alignment, addressing both the advancements of AI systems to incorporate human values, and the evolving human adaptation of AI systems in real-world practice.
4. Bidirectional Human-AI Alignment Framework
Figure 2. The fine-grained topology of “Align AI to Humans” direction, which studies mechanisms to ensure that AI system’s objectives match those of humans’. The goal is to integrate human specification to train, steer and customize AI systems.
This section introduces the Bidirectional Human-AI Alignment framework, which encompassing two interconnected alignment directions, as shown in Figure 1. The [Uncaptioned image] “Align AI to Humans” direction studies mechanisms to ensure that AI systems’ values match those of humans’ (Section 4.1 and 4.2). We identify two crucial research questions to be addressed in this direction: RQ1. What relevant human values are studied for AI alignment, and how do humans specify these values? and RQ2. How can human values be integrated into the development of AI? In comparison, the [Uncaptioned image] “Align Humans to AI” direction investigates the humans’ cognitive and behavioral adaptation to the AI advancement (Section 4.3 and 4.4). We recognize two imperative research questions along this direction: RQ3. How might humans learn to perceive, explain, and critique AI? and RQ4. How do humans and society make behavioral changes and react to AI advancement?. We address the four questions by elucidating the fine-grained typologies and literature solutions derived from our systematic review, as delineated in Figure 2 and Figure 3, respectively.
DIRECTION-I:  [Uncaptioned image] ALIGN AI to HUMANS
This direction delineates alignment research from AI-centered perspective (e.g., ML/NLP domains) and provides AI developers and researchers with approaches for two main challenges: carefully specifying the values of the system, and ensuring that system adopts the specification robustly (Wikipedia, 2024; Ngo et al., 2024). Therefore, as shown in Figure 2, we explore the two core research questions in this direction as: RQ1. Human Values and Specifications (Section 4.1) and RQ2. Integrating Human Specifications into AI (Section 4.2)
4.1. Align AI to Humans: [Uncaptioned image] Human Values and Specifications
RQ1: What relevant human values are studied for AI alignment, and how do humans specify these values? We structure existing literature to address this research question by answering the following “Sub-Research Questions”: What values have been aligned by AI? (Section 4.1.1) and then exploring How humans could interactively specify values in AI development? (Section 4.1.2). As shown at the top of Figure 2, particularly, we articulate the “Dimensions“ we summarized to answer each of these sub-research questions, and provide “Codes“ associated with “Example Papers” that have been studied in each dimension.
4.1.1. Categorizing Aligned Human Values. What values have been aligned with AI?
To clarify and systematically understand human values relevant to human-AI alignment, we leverage the adapted “Schwartz Theory of Basic Values” introduced in Section 3.3, and examine the category of aligned human values from the two dimensions of “Sources” and “Types”.
Sources of Values. This dimension examines the three sources of human values (Schwartz, 1994). Individual sources indicate values from individuals comprise universal needs of individuals as biological organisms (Schwartz, 1994) or prioritize personal interests (Kiesel et al., 2022; Jin et al., 2022). At this level, value alignment can be usually assessed independently of the context of interaction. This perspective highlights the values about technical capabilities of AI models, including factuality (Roit et al., 2023; Hu et al., 2024), calibration (Liu et al., 2024a), output diversity (Bradley et al., 2024), and model inductive bias (Si et al., 2023). Additionally, this code includes research on aligning model behaviors with the characteristics and preferences of individual humans, e.g., predict human moral judgements and decisions (Jin et al., 2022), and cognitive biases (Jones and Steinhardt, 2022; Koo et al., 2023). Social sources mean values from the social groups include universal requirements for smooth functioning and survival of groups (Schwartz, 1994; Ammanabrolu et al., 2022). Value alignment at this level transcends personal interactions and emphasizes the broader categories defined by shared experiences, identities, cultures, norms, and more (Santy et al., 2023). Research in this area often targets the alignment of AI behaviors with the general preferences of humans (Hendrycks et al., 2021; Bakker et al., 2022; Lee et al., 2023a). Additionally, there are efforts aimed to align AI with specific targeted groups, examining issues through the lenses of fairness (Fleisig et al., 2023), social norms (Sky et al., 2023), morality (Prabhumoye et al., 2021; Rao et al., 2023), and beyond. Interactive sources considers values at the interaction level to include universal requisites of coordinated social interaction (Schwartz, 1994, 2012), which typically occurs in interpersonal situations, such as in the dynamics of speaker-receiver relationships during language communication (Hovy and Yang, 2021; Cabrera et al., 2023). In the context of “human-AI alignment”, we adjust the definition of Interaction values to be the AI values that humans expect for AI as tools, such as Usability (Zhang et al., 2023a), Autonomy (Wang et al., 2023a), among others. Research along this line focuses on alignment strategies to enhance human-AI interaction (Gupta et al., 2023; Oh et al., 2023), collaborative decision-making (Gajos and Mamykina, 2022; Yang et al., 2023a), and trust (Buçinca et al., 2021; Dhuliawala et al., 2023).
Types of Values. In this dimension, we introduce the five high-order human values derived from a combination of the Schwartz Human Values (Schwartz, 1994) and empirical values studies from the surveyed papers. We provide a more in-depth taxonomy of value types in Section 5.2, including the relationship of these five types and more nuanced value categories. Self-Enhancement refers to a set of self-protective and personal values that emphasize enhancing self-esteem and a sense of personal worth (Sedikides and Strube, 1995; He et al., 2023; Xu et al., 2024; Huang et al., 2023; Park et al., 2023; Kang et al., 2023; Ye et al., 2024; Lima et al., 2021; Santurkar et al., 2023). One of the most important aspect is achievement, is competence as judged by social standards, which includes general capability (effectiveness/efficiency) as covered by the majority of the research (Guo et al., 2024). Another dimension is power, which relates to social status and prestige, as well as control or dominance over people and resources (Pinski et al., 2023). Self-Transcendence refers to a set of self-expanding and socially-focused human values that emphasize expanding beyond oneself (Frankl, 1966; Banovic et al., 2023; Wu et al., 2023d; Jin et al., 2022; Wu et al., 2023b; Huang et al., 2023; Ma et al., 2023b; Isajanyan et al., 2024; He et al., 2023; Nie et al., 2023; Santurkar et al., 2023; Ramezani and Xu, 2023; Sheng et al., 2021; Goldfarb-Tarrant et al., 2023). One critical aspect is benevolence, which relates to preservation and enhancement of the welfare of people with whom one is in frequent personal contact. Extensive efforts have investigated the values including helpfulness (Bai et al., 2022a), honesty/factuality (Hu et al., 2024), responsibility/accountability (Solyst et al., 2023). Another critical aspect is universalism, which relates to understanding, appreciation, tolerance and protection for the welfare of all people and for nature. Researchers have looked into values including inclusion/broad-mindedness (Bakker et al., 2022; Dong et al., 2024) and equality/fairness (Shen et al., 2022; Wang et al., 2023b). Conservation refers to a set of self-protective and socially-focused human values that hold and safeguard traditional institutions and customs (Hamilton, 2020; Qi et al., 2023; Zagalsky et al., 2021; Lai et al., 2022; Santy et al., 2023; Zhu et al., 2023; Zhou et al., 2023a). Under conservative values, people are concerned about security (Ssafety, harmony, and stability of society, relationships, and oneself), tradition (Respect for and acceptance of the customs and ideas provided by traditional culture or religion), and conformity (restraint of actions, inclinations, and impulses likely to upset or harm others and violate social expectations or norms). Exemplar studies in this field include safety (Anwar et al., 2024), mental health (Liu et al., 2023), and cultural moral norms (Ramezani and Xu, 2023). Openness to Change refers to a set of self-expanding and personally-focused human values motivated by an anxiety-free need to grow, in contrast to conservation (Bakker et al., 2022; Lima et al., 2020; Ashktorab et al., 2020; Jiang et al., 2023; Lima et al., 2020; Kang et al., 2023; Varanasi and Goyal, 2023; Yan et al., 2023; Hosking et al., 2024; Wu et al., 2023d; Kapania et al., 2023). Under this type, people are concerned about stimulation (seeking excitement, novelty, and challenges in life), hedonism (pursuing pleasure and sensuous gratification for oneself), and self-direction (independent thought and action—choices, creativity, and exploration). Exemplar studies in this field include understanding of privacy (Mireshghallah et al., 2024) and creativity (Isajanyan et al., 2024; Ashkinaze et al., 2024). We adopt the four high-order human values from Schwartz Human Values (Schwartz, 1994), supplementing them with additional values empirically collected from survey papers as shown in Figure 2. Additionally, as the conventional theory is missing the context of AI-system and human-AI interactions, we introduce a new high-order value type called Desired Values for AI Tools. This category encompasses the values humans expect from AI when used as tools in “human-AI interaction” (Yang et al., 2023b; Zhang et al., 2023a; Oh et al., 2023; Bakker et al., 2022; Boyd, 2021; Wang et al., 2023a; Paleja et al., 2021). These values include usability (competence according to the human experience on AI functionality) and human-likeness (resemblance to human intelligence and behavior). Exemplary studies in this area include assessments of AI usability (Zhang et al., 2023a) and autonomy (Wang et al., 2023a).
4.1.2. Interaction Techniques to Specify AI Values. How humans could interactively specify values in AI development?
This sub-research question investigates how human values are interactively111 We define ”interactions” broadly to encompass both ”synchronous” and ”asynchronous” interactions between humans and AI: (1) “Synchronous Interactions” indicates real-time exchanges where humans and AI systems interact simultaneously. (2) “Asynchronous Interactions” allows for delays between actions and responses, such as data annotations by humans to train the AI models. specified for AI systems to ensure alignment. It aims to elucidate the interaction techniques by which AI systems manifest or instantiate human values, thereby revealing the underlying mechanisms that shape their behavior or functionality.
Explicit Human Feedback. This dimension refers to the direct specification of human values through explicitly defined formats or mechanisms. Principles provides AI systems with explicitly defined principles, guidelines, or rules that dictate behavior or decision-making in alignment with human values (Petridis et al., 2024a; Bai et al., 2022b). Rating and Ranking is widely used to assign numerical scores or rankings to options or outcomes based on their alignment with human values (Dai et al., 2024; Yuan et al., 2023). Natural Language Interaction/Conversations allows humans to interact with AI systems through natural language interfaces to express and communicate human values (Bai et al., 2022b; Hwang et al., 2023; Guo et al., 2024; Feng and Boyd-Graber, 2019; Yao et al., 2023a; Lin et al., 2024; Hong et al., 2022; Shen et al., 2023a; Zhang et al., 2021). For Multimodal Feedback, human values can also be provided in multiple modalities, such as sketches/images and gestures (Chung et al., 2022), to convey human values (Chen et al., 2022; Geva et al., 2022; Wu et al., 2022; Jörke et al., 2023).
Implicit Human Feedback. This dimension refers to the indirect representation or inference of human values within AI systems through patterns, signals, or cues embedded in the data or decision-making processes. Discarded Options refers to the options or choices that human discard when interacting with AI systems during the decision-making processes, which also potentially infer human values (Park et al., 2021b). Language Analysis means that textual data and language patterns can also contain rich information to identify implicit references to human values or value-related concepts (Lu et al., 2023). Theory of Mind refers to the ability of agents and people to attribute mental states, such as beliefs, intentions, desires, emotions, knowledge, percepts, and non-literal communication, to themselves and others (Ma et al., 2023b). Social Relationships refers to the implicit values derived from analyzing human social relations and behaviors, which are derived from external sources (e.g., social network) that can inherently reflect their values (Isajanyan et al., 2024).
Simulated Human Value Feedback. When human values in explicit formats are expensive or impossible to collect, one may simulate human-like feedback within AI systems to approximate human responses and preferences regarding specific values. Human Feedback Simulators uses computational algorithms to simulate human-like feedback on values, based on predefined criteria or training data (Petridis et al., 2024b; Li et al., 2023a). Comparison to Human Data refers to developing techniques that assess the likelihood or probability of AI-generated outputs matching human behaviors in a reference set or dataset (Dong et al., 2024). Synthetic Data curates data by generating synthetic comparisons based on naive assumptions or heuristic rules, followed by post-validation to ensure feedback quality (Kim et al., 2023).
4.2. Align AI to Humans: [Uncaptioned image] Integrating Human Specifications into AI
RQ2. How can human values be integrated into the development of AI? Existing studies have explored diverse methods to integrate human values into AI. We structure them by summarizing how to integrate general human values (Section 4.2.1) and customized human values (Section 4.2.2) throughout AI development stages?, and then elaborating what are the evaluation methods (Section 4.2.3) and supported platforms (Section 4.2.4) for the AI development? Additionally, we answer the four “Sub-Research Questions” by introducing the answer “Dimensions“ and providing “Codes“ associated with “Example Papers”.
4.2.1. Integrating General Values to AI: how to incorporate general human values into AI development?
This sub-research question focuses on the process of incorporating broad, universally recognized human values into the development of AI systems. The goal is to ensure that AI systems align with overarching ethical principles and societal norms, thereby promoting trust, acceptance, and responsible use.
Instruction Data. This dimension refers to the types of data and processes used to provide guidance or direction to AI systems during their development. Human Annotation makes data with human-generated labels or annotations that indicate the presence or relevance of specific human values (Shen et al., 2023b; Pei et al., 2022). Human-AI CoAnnotation leverages both human expertise and AI capabilities to collaboratively annotate the data  (Sun et al., 2023; Li et al., 2023b). Simulated Human Data generates synthetic or simulated data that mimics human behaviors, preferences, or decision-making processes to provide training signals for AI systems (Dubois et al., 2023; Kim et al., 2023).
Model Learning. This dimension refers to the model architecture design and training stages after the data collection, where human values are integrated during the model learning process. Online Human Alignment integrates human values into AI systems in real-time or during active system operation, often through interactive feedback loops or adaptive learning mechanisms. Examples include real-time user feedback and online training (Ouyang et al., 2022; Dong et al., 2023). Offline Human Alignment incorporates human values into AI systems prior to deployment or during offline training phases, without direct user interaction (Rafailov et al., 2023; Song et al., 2024; Yuan et al., 2023).
Inference Stage. This dimension involves evaluating the alignment of AI systems with human values and assessing their performance and behavior in relation to predefined criteria or benchmarks. Prompting leverages prompting methods, such as in-context learning and chain-of-thought, on trained AI systems to elicit or critique AI regarding the encoded values (Bai et al., 2022b; Wang et al., 2023a). External Tool Interaction integrates external tools, like code interpreter for debugging, to cross-check and refine their initial generated content (Gou et al., 2024). Response Search generates a diverse range of high-quality outputs from which to choose (Bradley et al., 2024).
4.2.2. Customizing AI Values: how to customize AI to incorporate values from individuals or human groups?
This sub-research question explores the customization of AI values to align with specific contexts, domains, or user preferences. The goal is to enhance the alignment of AI systems within specific application domains or user communities.
Customized Data. Finetuning with Curated Datasets aims to curate datasets for specific individuals or societal groups, and further finetune the pre-trained AI models on these specific datasetsto align them with targeted human groups and values (Felkner et al., 2023). These curated datasets include data collected from socio-demographic groups (Orlikowski et al., 2023), users’ history data (Lyu et al., 2023), expert-selected data for imitation learning (Cai et al., 2023) and others.
Adapt Model by Learning. This dimension involves refining or adjusting AI values through techniques for customization, such as iterative learning process, model enhancements, or structural modifications. Group-based Learning trains AI models from specific user groups or communities to capture group-specific values or preferences (Zhao et al., 2023). Active Learning aims to interative selecting and lebeling data samples for AI model training based on their potential to improve alignment with user preference or values (Peschl et al., 2022). Inserted Adapters incorporates adapter modules or components into AI model architectures to fine-tune specific aspects or behaviors (Sun et al., 2022). Mixture of Experts combines multiple specialized models or experts to collectively capture diverse perspectives and values, with each expert focusing on a specific subset of the data or problem space (Shen et al., 2022). Enhanced Knowledge enhances AI model’s representations and embeddings with additional knowledge or context to improve alignment with user preferences or values (Wang et al., 2023c; Dalvi et al., 2022).
Interactive Alignment This dimension involves actively engaging users or stakeholders in the process of customizing AI values to align with specific contexts, domains, or user preferences. Interactive Learning enables the users to provide feedback or corrections to AI models in real time, such as using interactive tutorials and user-driven customization interfaces (Maghakian et al., 2023; Swazinna et al., 2023). Steering Prompts provides users with prompts or cues to steer the behavior or decision-making of AI systems towards desired outcomes or values (Lahoti et al., 2023; Fei et al., 2023). Proactive Alignment anticipates user needs and preferences based on historical data or user profiles and proactively adjusting AI systems accordingly (Welch et al., 2022; Shen et al., 2023a).
4.2.3. Evaluating AI Systems: how to evaluate AI regarding human values?
The rise in the use of LLMs has also seen the rise of automatic evaluation of generated natural language text evaluation in different contexts. But in particular, a research dimension has focused on analyzing how closely values discussed in humane context has been adapted to AI models/applications and how these are being evaluated.
Human-In-The-Loop Evaluation. This dimension involves incorporating human judgement, feedback, or interaction into the evaluation process to assess the effectiveness, robustness, and ethical implications of integrating human values into AI systems. Human Evaluation solicits feedback, opinions, or assessments from human evaluators to gauge the alignment of AI systems with human values and ethical standards (Liu et al., 2022; Sharma et al., 2023; Ouyang et al., 2023; Kiesel et al., 2022; Hosking et al., 2024). Human-AI Collaborative Evaluation collaboratively evaluates AI systems with both human evaluators and large language models (LLMs) to leverage the strengths of both human judgement and AI capabilities (Kwon and Mihindukulasooriya, 2023; Ammanabrolu et al., 2022).
Automatic Evaluation. This dimension involves using computational methods or algorithms to assess the alignment of AI systems with human values, without direct human involvement. Human Simulators use simulation models or virtual agents to mimic human behavior and assess AI performance in human-like scenarios. Typical methods include agent-based simulations, synthetic user models, and others (Jia et al., 2024; Wang et al., 2023d; Felkner et al., 2023; Ruan et al., 2023). Evaluation Benchmarks establishes standardized benchmarks or metrics for evaluating AI performance in relation to human values and ethical considerations (Sun et al., 2023; Ramezani and Xu, 2023; Hu et al., 2024; Kiesel et al., 2022; Rao et al., 2023; Lee et al., 2023a; Sheng et al., 2021; Sharma et al., 2023). Distribution Difference compares the difference between the output distribution from AI generations and human data to evaluate AI (Patel and Pavlick, 2021; Lee et al., 2023a).
Figure 3. The fine-grained topology of “Align Human to AI” direction, which studies humans’ cognitive and behavioral adaptation to the AI systems, which aims to help humans better critique, collaborate with, and co-adapt to AI.
4.2.4. Ecosystem and Platforms: how to build the ecosystem to facilitate human-AI alignment?
The ecosystem and platforms refer to the broader context in which AI systems operate and interact with other agents, platforms, or environments. This includes the infrastructure, frameworks, and technologies that support the development, deployment, and utilization of AI systems. LLM-based Agents are based on large language models (LLMs) such as GPT (Generative Pre-trained Transformer) models, which have been pre-trained on vast amounts of text data (Dubois et al., 2023; Xu et al., 2023; Wu et al., 2023a; Zhou et al., 2023b). RL-based Agents are based on reinforcement learning (RL) algorithms to learn and adapt their behavior based on feedback from the environment or human users (Yuan et al., 2024). Annotation Platforms refers to the ecosystems that are designed to crowdsource human demonstrations as collected data for reinforcement learning (Gerstgrasser et al., 2021) and supervised finetuning learning for alignment (Pei et al., 2022).
    Underexplored Dimensions in Aligning AI to Humans (See Sec 5.1 for supporting data and evidence): 1. Implicit human feedback and simulated human value feedback are under-explored in existing research work. 2. Developing and customizing AI during the inference stage or in an interactive way is under-explored. 3. Human-in-the-loop evaluation is much less explored than automatic evaluation.
DIRECTION-II: [Uncaptioned image] ALIGN HUMANS to AI
[…]
When interacting with people, AI agents do not just influence the state of the world – they also influence the actions people take in response to the agent, and even their underlying intentions and strategies (Hong et al., 2023). From a long-term perspective, it is essential to consider the dynamic changes around human-AI alignment. This direction outlines alignment research from the Human-centered perspective (e.g., HCI/Social Science domains) and provides HCI researchers and user experience designers with guidance for approaching two core research questions: RQ3. Human Cognitive Adjustment to AI (Section 4.3) and RQ4. Human Adaptive Behavior to AI (Section 4.4).
4.3. Align Humans to AI: [Uncaptioned image] Human’s Perceptual Adaptation to AI
RQ3. How might humans learn to perceive, explain, and critique AI? Humans needs to understand AI to better specify their demands and collaborate with AI. Also, as AI systems produce a range of risks, it is important to elicit humans’ critical thinking of AI instead of relying on AI blindly. Therefore, we categorize existing literature to answer the questions of how humans learn to perceive and understand AI (Section 4.3.1), and how to engage in critical thinking on AI? (Section 4.3.2). We also answer the two “Sub-Research Questions” by introducing the answer “Dimensions“ and providing “Codes“ associated with “Example Papers”.
4.3.1. Perceiving and Understanding AI: how do humans learn to perceive and explain AI systems?
Addressing the problem of human perception and understanding of AI includes the fundamental education and training required for less technical people to improve understanding of the behind mechanisms and outputs produced by AI systems. It includes visualizations and human-centered explanations tohelp more people learn to understand AI outputs.
Educating and Training Humans. AI is increasingly incorporated into the daily lives of a broad spectrum of users, including those with little to no technical knowledge on how AI operates. Here we discuss how to help these less technical people become more AI literate, or better trained to use AI. AI Literacy and Awareness is broadly defined as the core competencies required for less technical people to better use and collaborate with AI (Long and Magerko, 2020). Beyond, there is also more in depth and explicit Training Courses to support individuals to better collaborate and utilize AI (Paranjape et al., 2019; McDonald and Pan, 2020).
AI Sensemaking and Explanations. To improve people’s understanding of AI systems and generated outputs, various techniques and approaches have been developed to help people learn to make sense of and explain the model’s outputs. Perception and Sensemaking on AI involves the process through which humans learn to make better sense of AI mechanisms and decision makings (Kaur et al., 2022; Sivaraman et al., 2023; Petridis et al., 2023). People also need to understand how AI systems arrive at specific generated outputs. Prior studies in Human-Centered Explanations have examined various approaches and interactive techniques to increase human understanding of AI generations and outputs (Tenney et al., 2020; Tintarev and Masthoff, 2007; Smith-Renner et al., 2020).
4.3.2. Critical Thinking around AI: how do humans think critically about AI systems?
Beyond simply perceiving and understanding AI, individuals need to compare their mental model of AI with their own mental model to judge whether the AI is behaving rationally and ethically. Broadly, this involves exploring the ways in which humans engage in critical thinking and reflect on their interactions with and evolving understanding of AI technologies. Here we discuss studies that help humans become more capable to identify biases and errors in AI output, and the ethical implications that arise from using AI algorithms in decision-making processes. Also, we discuss how humans need to calibrate their mental models to be more aligned with AI.
Trust and Reliance on AI Decisions refers to the extent to which people mentally trust the AI competency and practically rely on AI for output generation and decision making. We use the term Trustworthiness to indicate whether humans decide to trust the reliability, integrity, and competence of an AI system in delivering accurate and trustworthy decisions or recommendations to the users (Yin et al., 2019; Ma et al., 2023a; Srivastava et al., 2023). We define AI Reliability as to what extent humans utilize and rely on AI to automate decision making in practice with low error rates and robustness performance (Schemmer et al., 2023). (Schemmer et al., 2023; Vasconcelos et al., 2023; Suresh et al., 2022). Selective AI Adoption refers to the criteria and considerations that guide the adoption or rejection of AI technologies based on their perceived benefits, risks, and alignment with user needs and preferences (Gu et al., 2024; Vasconcelos et al., 2023).
Ethical Considerations and AI Auditing refers to potential moral and societal issues that arise from the development, deployment and use of AI systems, as well as systematic examination and evaluation of AI systems regarding these issues. Ethical Concerns and AI Mitigation encourages humans to carefully consider if the AI systems possess ethical concerns (e.g., such as bias and discrimination, privacy violations and the potential for harm or misuse) and implement strategies and practices to address and reduce these ethical concerns associated with AI (Holstein et al., 2019). Audit Deployed AI refers to the systematic examination of AI systems to ensure that they operate as intended, comply with ethical and legal standards, and do not cause unintended harm (Bandy, 2021).
Re-Calibrating Cognition to Align with AI. This sub-category deals with interventions and techniques to help users adjust (1) their own mental model of how the AI operates and (2) recalibrate their perception of the AI. Updating Human Mental Model of AI means that when the AI is not aligned with humans, it is important to adjust human perceptions of the confidence in AI systems based on the model’s capability and track record (He et al., 2023; Kocielnik et al., 2019) . Recalibrating Trust and Reliability indicates that humans adjust their perceptions of trust and reliability in AI systems based on their performance and reliability. This is also important to foster appropriate reliance and skepticism (Wischnewski et al., 2023).
4.4. Align Humans to AI: [Uncaptioned image] Human’s Behavioral Adaptation to AI
RQ4. How do humans and society make behavioral changes and react to AI advancement? As AI becomes increasingly integrated into daily life, it is essential to understand its influence on humans, encompassing both positive and negative aspects. Moreover, it is crucial to determine how individuals and society can best and most appropriately respond to this influence. To this end, we summarize literature to answer how do humans learn to collaborate with AI in diverse AI roles? (Section 4.4.1), how humans and society are impacted by AI (Section 4.4.2) and how might we assess these impacts? (Section 4.4.3) We articulate the three “Sub-Research Questions” by introducing the answer “Dimensions“ and providing “Codes“ associated with “Example Papers”.
4.4.1. Human-AI Collaboration Mechanisms: what are human strategies to collaborate with AI that have differing levels of capabilities?
This category looks at many ways that humans and AI can collaborate, such as teamwork, co-creation, and coproduction.
AI Assistant for Humans captures the essence of a symbiotic relationship where AI systems are designed to bolster human capabilities, with humans steering the interactions. LLM-based Demand Specification employs LLMs to interpret and respond to human requests, powering virtual assistants and chatbots that streamline information retrieval and improve task accuracy (Terry et al., 2023). This naturally extends to LLM-based Prompt Strategy, which helps humans to better write prompts using additional abstraction and scaffolding methods. The example system can enhance humans’ capability in generating intelligent prompts and suggestions, such as autocomplete and question-generation tools for LLM-based systems (Wu et al., 2022; Wang et al., 2024), facilitating humans’ smooth and intuitive decision-making processes. In the creative arena, LLM-based Prototypes utilize AI to transform human ideas into tangible or organized concepts  (Gebreegziabher et al., 2023), enabling professionals to explore and refine a wide array of artistic possibilities with AI-generated options (Huh et al., 2023). On a broader note, researchers have also explored how AI-Assisted Task and Decisions aims to achieve complementary performance for human-AI collaborative tasks by empower humans to discern when and how to adopt AI-assisted recommendations or AI-generated explanations for decision-making (Choi et al., 2019; Wang et al., 2022; Wang and Yin, 2023; Zheng et al., 2023; Wu et al., 2023c).
Human-AI Partnership refers to a collaborative relationship where humans and AI systems work together as partners, combining their respective strengths to achieve shared goals more effectively. Simulated Agency enables humans to collaborate with AI partners with simulated agency or autonomy (e.g., autonomous agents and collaborative robots (Park et al., 2023)) to make decisions collaboratively. Reciprocal Learning focuses on how humans learn from and exchange knowledge with AI systems, enhancing human-AI collective capabilities and performance through knowledge-sharing platforms and collaborative filtering (Zagalsky et al., 2021; Pinski et al., 2023). AI Delegation allows humans to delegate AI partners to help finish tasks or responsibilities, facilitated by task assignment algorithms and workflow management systems (Pinski et al., 2023). Furthermore, AI-Mediated Tasks emphasizes how traditional human tasks or behaviors (e.g., communication) would be changed by incorporating AI partners in the loop (Bernstein et al., 2010; Sundar and Lee, 2022; Xu and Zhang, 2023). Co-Design with AI explores AI as a design material or a partner in prototyping, which enables humans to converse with AI in situations that can collaboratively improve the design outcomes. (Yildirim et al., 2022; Wang et al., 2024)
AI Tutoring for Human Learning investigates how humans improve their learning and knowledge throughinteractions with AI tutors that can perform better than humans in some tasks. With the tailored instructions and customized feedback from AI tutors, humans can enhance their the learning outcomes and facilitate mastery of new skills more effectively than traditional learning methods. AI Tutor for Technical Skills refers to empowering humans to learn technical skills from AI tutors. For technical subjects like coding, AI tutors can analyze a learner’s progress in real time, adjusting the pace, content, and approach to ensure a solid understanding of complex concepts and practical skills such as programming (Ma et al., 2024). Similarly, AI Tutor for Social and Behavioral Skills involves enabling humans to learn social and behavioral skills from AI tutors. Humans can leverage virtual simulations, created by AI tutoring systems, to practice public speaking, interpersonal communication, and other soft skills. By analyzing verbal and non-verbal cues, humans can receive constructive feedback on areas such as body language, tone, and delivery from AI, ultimately enhancing their ability to communicate effectively across various settings (Peng et al., 2021; Shaikh et al., 2024).
4.4.2. AI Impact on Humans and Society: how are humans influenced by AI systems ?
This category explores the effects of AI advancement on human behaviors, attitudes, and societal dynamics. It involves examining the behavioral changes, adaptations, and reactions that individuals, groups and wider communities undergo in response to the proliferation of AI technologies. The goal is to elucidate the multifaceted impacts of AI on human behavior and society and to inform policy-making, education, and intervention efforts.
Impacts on Participatory Individuals and Groups covers the effects of AI advancement on the behaviors, attitudes, and experiences of both individuals and groups. This dimension focuses on examining how AI technologies influence decision-making, creativity, privacy, and authorship. Decision Making refers to analyzing how AI technologies influence human decision-making processes, including biases, preferences, and risk assessment, in various domains such as healthcare, finance, and personal life (Shen and Huang, 2020). Human Creativity explores the impact of AI technologies on human creativity, innovation, and expression, including the augmentation or automation of creative tasks and the emergence of new forms of artistic expression (Ashkinaze et al., 2024). Privacy relates to investigating the implications of AI technologies for individual privacy rights, data protection, and surveillance, including concerns about data collection, tracking, and algorithmic profiling (Li et al., 2024a). Authorship catalogs issues related to intellectual property, attribution, and ownership of AI-generated content, including questions of legal responsibility, copyright infringement, and plagiarism detection. Salient is that AI can produce increasingly realistic, synthetic data quickly and at low cost, which brings forth tensions around the use of such data to make decisions (Hämäläinen et al., 2023; Atreja et al., 2023).
Societal Concerns and AI Impacts involves the broader societal implications and consequences of AI advancement on misinformation, education, social relationships, norms, job displacement, and other aspects of human society. Misinformation and Moderation concerns the challenges of misinformation, disinformation, and online content moderation in the context of AI-driven information ecosystems, including concerns about algorithmic bias and filter bubbles (Atreja et al., 2023). Impacts on Education pertains to assessing the effects of AI technologies on education systems, learning outcomes, pedagogical practices, and workforce training, including opportunities for personalized learning and skill development (Kazemitabaar et al., 2023). Impacts on Social Relationship and Norms explores how AI technologies shape interpersonal relationships, social interactions, and cultural norms, including changes in communication patterns, social dynamics, and ethical considerations (Cheng et al., 2022). Workplace refers to examining the effects of automation and AI technologies on employment patterns, job markets, and workforce dynamics, including concerns about job displacement, re-skilling, and economic inequality (Park et al., 2021a).
Reaction to AI Advancement involves societal responses, regulatory frameworks, and policy initiatives aimed at addressing the challenges and opportunities posed by AI technologies. This dimension encompasses efforts to regulate AI deployment, re-calibrate societal acceptance, and manage potential backlash. For example, reaction to bias and discrimination in algorithmic decision making can depend on how people perceive the machine and the context of use, i.e., if the machine is considered an actor embedded in social structures that call for blame when harmful decisions are made (Lima et al., 2023a). AI Regulatory and Policy includes regulatory frameworks, legal frameworks, and policy initiatives aimed at governing AI development, deployment, and use, including concerns about ethics, safety, and accountability (Hacker et al., 2023; Lucaj et al., 2023). Shifts in AI Acceptance relates to investigating societal attitudes, organizational practices, and acceptance of AI technologies over time in practice, including shifts in public opinion and AI utilization by humans and institutes regarding AI deployment and impact (Lima et al., 2023b; Pruss, 2023). Surveillance and Intervention emphasizes the need for transparency, monitoring, and human oversight in the algorithmic decision-making process. This approach enables better human control over AI systems and helps mitigate potential risks associated with their use (Park et al., 2021a). Responsible AI Checklists involves the creation of ethical guidelines, such as those focusing on fairness and transparency, to ensure the responsible development and deployment of AI systems. These published principles serve as a foundation for guiding ethical AI practices (Madaio et al., 2020).
4.4.3. Evaluation in Human Studies: how might we evaluate and understand the impact of AI on humans and society?
We summarize common empirical methods used to rigorously understand and assess the impact of AI on humans. Specifically, we focus on two types of impact. On the micro-level, we discuss how to evaluate the effectiveness of human-AI collaboration; on the macro-level, we discuss how to assess the impact of AI on a large group of people over a long period of time.
Evaluate Human-AI Collaboration refers to the evaluation of the effectiveness of an AI system in collaboration with humans. It is key to not only consider the final output, but also the interaction experience (Lee et al., 2023b). Human-AI Team Performance compares human-AI team performance with the performance of humans alone without AI collaboration The metrics for measuring performance should include both task success metrics (e.g., accuracy) as well as indicators of efficiency (Mozannar et al., 2024; Ding et al., 2023a). Cognitive Workload and User Satisfaction involves understanding the degree of cognitive load that the user experiences when interacting with the system, as well as user satisfaction with the interaction and the final outcome. Such aspects are often captured via surveys or interviews (Ding et al., 2023a; Lee et al., 2023b). Control and Trust refers to how user control can support the avoidance of catastrophic AI failures, especially in high-stakes settings where AI mistakes could lead to harm (Srikanth et al., 2024; Ding et al., 2023b). Interaction Behavioral Analytics refers to measuring task performance quantitatively. This approach includes recording user interaction data and analyzing the patterns (Villalba et al., 2023; Perry et al., 2023). Qualitative Survey and Interviews refers to qualitative approaches to understanding human-AI interaction. Commonly used methods include qualitative survey questions (i.e., open-ended) and user interviews to assess aspects of the user experience. (Zhang et al., 2023b). Statistical Analysis utilizes methods such as regression analysis to quantitatively analyze and evaluate data from human studies, allowing for the verification of hypotheses (Wang et al., 2021).
Evaluate Societal Impact refers to the macro-impact of a group of people as they come to use AI broadly. This dimension requires sufficient scale and time. The aim is to understand how the group’s behavior changes as people within it frequently interact with AI. Public Opinion Surveys aims to investigate the impact of AI on human measures of interest though deploying and analyzing large scale questionnaires  (Santurkar et al., 2023). Behavioral Data Analytics collects large-scale and potentially longitudinal behavior data, with the aim of understanding how pattern evolve over time (Shin et al., 2023; Zhou and Lee, 2024).
    Underexplored Dimensions in Aligning Humans to AI (See Sec 5.1 for supporting data and evidence): 1. Educating and training humans on AI literacy is under-investigated. 2. Auditing AI for various ethical values is not fully explored. 3. The collaboration between humans and AI with similar or superior capabilities is under-explored. 4. The societal impacts of and reactions to AI advancements are not fully explored.
5. Findings and Discussions on Current Gaps
In this section, our aim is to consolidate key findings derived from our analysis of the framework and the current state of literature we reviewed. We begin by analyzing the overall trends and gaps in the literature (Section 5.1). We then focus on three essential aspects: the relationship between human values and alignment (Section 5.2), the potential interaction modes used to specify human values (Section 5.3).
Figure 4. The number of papers for each dimension in the bidirectional human-AI alignment framework. Out of papers that are relevant to each research question (i.e., gray bars), we show the number of papers that are relevant to each dimension (i.e., color bars). This figure illustrates the extent to which each dimension has been explored by existing research. We provide more analysis in Section 5.1.
5.1. Meta Analysis of Trends and Gaps
Based on our coding of all papers, we computed the number of relevant papers for each dimension in the proposed bi-directional framework. Based on the distribution in Figure 4, we noticed that in existing literature certain dimensions are over- or under-represented. We outline the under-explored dimensions in two directions in the highlighted grey bars in the end of Section 4.2 and Section 4.4, respectively, and provide more details below.
Underexplored dimensions in AI-centered alignment research. Most literature specified human values using explicit human feedback, whereas implicit and simulated human feedback were largely under-explored. In the integrating human specifications to AI research question, Besides, developing and customizing AI models during the inference stage or using the interactive approach is under-explored. Also, human-in-the-loop evaluation is much less explored than automatic evaluation.
Underexplored dimensions in Human-centered alignment research. Additionally, many studies on human cognitive adjustment to AI focus on enabling AI sensemaking and explanations so that humans can better understand, trust, and rely on AI. However, these studies often focused on explaining AI decision-making justification rather than educating people to acquire general skills and competencies required to understand, use, critique, and interact effectively with AI systems (i.e., AI literacy). AI literacy (Long and Magerko, 2020) plays a fundamental role in ensuring people understand and use AI correctly, and warrants deeper exploration. Also, we observed a wide range of studies (Wu et al., 2022; Gebreegziabher et al., 2023; Park et al., 2023; Ma et al., 2024) which developed interactive mechanisms and prototypes to empower human collaboration with AI. However, most studies assumed that AI plays an assistant role, being less capable than humans. This situation might change over time. Moreover, the influence of AI advancements on human behavior, social relationships, and societal changes is essential but remains largely unexplored.
Figure 5. The interaction techniques for specifying values in human-AI alignment. We compare the common interaction techniques used for the model “Learning” and “Inference” stages in human-focused (e.g., HCI) and AI-focused (e.g., NLP/ML) research studies.
    Takeaways of Interaction Techniques for Alignment. 1. Some common human feedback formats (rating, ranking) used in NLP/ML are not often studied in HCI. 2. Diverse human interactive feedbacks in HCI are not fully used in AI development in NLP/ML fields.
5.2. Insights into Human Values for Alignment
Our analysis, based on the adaptation of Schwartz’s Theory of Basic Values and our comprehensive literature review, identifies three critical findings for future research:
[…]
Universal vs. Personalized AI Values. While certain values are universally expected from AI (e.g., capability, equity, responsibility), others may be undesirable in specific contexts (Ngo et al., 2024) (e.g., seeking power). Simultaneously, AI models should be adaptable to diverse human value systems (Sorensen et al., 2024). Research is needed to develop methods for identifying appropriate value sets for specific individuals or groups, and for customizing AI to align with user values while maintaining ethical principles.
Disparities in Value Expectations and Evaluation. The fundamental differences between humans and AI necessitate distinct approaches to value evaluation. For instance, assessing AI honesty may require mechanistic interpretability (Bereska and Gavves, 2024), a more rigorous standard than that applied to humans. Future studies should explore methods for evaluating and explaining AI values and calibrating human expectations accordingly.
5.3. Interaction Techniques for Specifying Human Values: A Cross-Domain Analysis
Our research reveals disparities in interaction techniques for human-AI value alignment across AI-centered (NLP/ML) and human-centered (HCI) domains. As depicted in Figure 5, this analysis focuses on three key areas:
Domain-Specific Interaction Techniques. The interaction techniques in AI-centered (NLP/ML) and Human-centered (HCI) alignment studies are often differ (Brath, 2021). NLP/ML studies primarily utilize numeric and natural language-based techniques. Also, NLP/ML research explore implicit feedback to extract human hidden feedback. In contrast, HCI research encompasses a broader range of graphical and multi-modal interaction signals (e.g., sketches, location information) beyond text and images. This disparity suggests potential gaps in extracting comprehensive human behavioral information.
Stage-Specific Interaction Techniques. In NLP/ML, the learning stage predominantly employs rating and ranking interactions for alignment in dataset generation. However, when humans use AI in the inference stage, as demonstrated in HCI research, involves more diverse user interactions. This discrepancy highlights the need for alignment between model development and practical deployment.
Divergent Data Utilization. NLP/ML typically uses interaction outputs as training datasets, while HCI analyzes this data to understand human behavior and feedback. As AI systems evolve, developing new interaction modes to capture a broader spectrum of human expression becomes crucial.
6. Future Directions
Drawing upon insights gained from the development of our framework and the associated coding analysis, as shown in Figure 6, we propose future research aiming to achieve the long-term alignment goal by identifying three important challenges from near-term to long-term objectives, including the Specification Game (Section 6.1), Dynamic Co-evolution of Alignment (Section 6.2), and Safeguarding Coadaptation (Section 6.3).
6.1. Specification Game
An important near-term challenge is resolving the “Specification Game”, which involves precisely defining and implementing AI goals and behaviors to align with human intentions and values. Next, we will introduce how synergistic efforts from two directions can potentially address this challenge.
[…]
Individuals often possess value systems that encompass multiple values with varying priorities, rather than a single value, to guide their behaviors (Schwartz, 1994, 2012). Also, these priorities can change dynamically throughout an individual’s life stages. As such, It is more realistic to select values compatible with specific societies or situations, given the fact that we live in a diverse world (Gabriel, 2020). Future research, inspired by Social Choice Theory (Arrow, 2012), could focus on using democratic processes to aggregate individual values into collective agreements. Building on the summaries in Sections  4.1 and  5.2, researchers can employ democratic methods to identify diverse subsets of human values for AI alignment. Additionally, creating datasets that represent these values is crucial. Besides, it is crucial yet challenging for AI designers to investigate how to fully specify the appropriate values and to further integrate these values into AI alignment. Future important area involves developing algorithms, such as the Bradley-Terry Model (Rafailov et al., 2023) or Elo Rating System (Bai et al., 2022b), to convert heterogeneous human values into AI-compatible formats for training reward models and guiding reinforcement learning. Researchers should also explore AI models capable of aligning with unstructured human data, including free-form descriptions of values, multimedia, or sensor recordings depicting human behavior.
[…]
Current alignment methods use instructions, ratings, and rankings to infer human values, which can not fully capture all relevant human values and constraints. Future research should focus on optimizing interactive interfaces to efficiently elicit human values. These interfaces can leverage diverse interaction modes to capture comprehensive human value information. Additionally, people often struggle to formulate optimal prompts for AI, accurately specify their requirements, and articulate their desired values, which can change based on context and time. Developing proactive interfaces that use conversational techniques to elicit nuanced and evolving values is also crucial. Implicit human signals that indicate values are also frequently overlooked. Additionally, systems that track interactions to hypothesize and validate implicit human values in real-time should be designed.
6.2. Dynamic Co-evolution of Alignment
The challenge ahead lies in comprehending and effectively navigating the dynamic interplay among human values, societal evolution, and the progression of AI technologies. Future studies in these directions aim to bolster a synergistic co-evolution between AI and human societies, adapting both to each other’s changes and advancements.
[…]
(i) Existing literature often treats AI alignment as static, ignoring its dynamic nature. A long-term perspective must consider the co-evolution of AI, humans, and society. As AI systems evolve and scale up, they gain new capabilities, making it essential to ensure their goals remain aligned with human values. Thus, alignment solutions require continuous oversight and updates. Future research should develop methods for continuously updating AI with limited data without compromising alignment values and performance. This could involve forecasting human value evolution and preparing AI with flexible strategies like prompting or interventions. (ii) Additionally, AI advancements also influence human actions and values, necessitating adaptive alignment solutions. Ensuring AI co-evolves with human and societal changes is crucial for robust alignment. This challenge could potentially be addressed by forecasting the potential evolution trajectories of human values or behavioral patterns, and preparing AI with the flexibility to adapt in advance, for example, through prompting or intervention strategies.
[…]
(i) While current AI systems lag behind humans in many tasks, identifying and handling AI mistakes, including knowing when to seek human intervention, remains essential. Future research should focus on developing validation mechanisms that enable humans to interpret and verify AI outputs. This could involve designing interfaces that allow humans to request step-by-step justifications from AI or integrating tools to verify the truthfulness of AI referring to Section 4.4.1. Additionally, developing interfaces that enable groups of humans to collaboratively validate AI outputs and creating scalable validation tools for large-scale applications are important directions. (ii) As AI advances, it becomes essential to develop systems that enable humans to utilize AI with capabilities surpassing their own. Research is needed to understand how individuals can interpret and validate AI outputs for tasks beyond their abilities and leverage advanced AI sustainably, avoiding issues like job displacement or loss of purpose. Another research direction is designing strategies to enhance human capabilities by learning from advanced AI, including gaining knowledge and building skills. (iii) As AI integrates more into daily tasks, its impact on human values, behaviors, capabilities, and society remains uncertain. Continuous examination of AI’s influence on individuals, social relationships, and broader societal changes is vital. Research should assess how humans and society adapt to AI advancements, guiding AI’s future evolution. Potential areas include evaluating changes in individual behavior, social relationships, and societal governance as AI replaces traditional human skills. Understanding these dynamic changes is essential for grasping the broader impact of AI on humanity and society.
6.3. Safeguarding Co-adaptation
As AI gains autonomy and capability, the risks associated with its instrumental actions, as a means toward accomplishing its final goals, increase. These actions can be undesirable for humans. Therefore, safeguarding the co-adaptation between humans and AI is crucial. We next explore future research to address this challenge from both directions.
[…]
(i) As advanced AI systems become more complex, they present greater challenges for human interpretation and control. It is crucial to empower humans to detect and interpret AI misconduct and enable human intervention to prevent power-seeking AI behavior. Research should focus on designing corrigible mechanisms for easy intervention and correction, including modular AI architectures and robust override protocols that allow human operators to halt or redirect AI activities. These components should be human-interpretable, enabling scenario testing. (ii) Furthermore, advanced AI systems may intentionally mislead or disobey humans, generating plausible fabrications (Johnson and Iziev, 2022). Developing reliable interpretability mechanisms to validate the faithfulness and honesty of AI behaviors is essential. This includes correlating AI behaviors with internal neuron activity signals, akin to physiological indicators in human polygraph tests (Association et al., 2004). Inspecting these indicators can help humans assess the truthfulness of AI interpretations and prevent risky actions.
[…]
(i) Preventing advanced AI from engaging in risky actions requires robust human supervision. Essential steps include developing training and simulation environments with scenario-based exercises and timely feedback, and creating interactive dashboards for real-time monitoring. These dashboards should feature effective data visualization, anomaly detection, and prompt alert systems for immediate intervention. (ii) Scalable solutions are needed for supervising AI across various applications. Real-time oversight becomes more challenging with widespread AI deployment, necessitating advanced autonomous monitoring tools. These tools should learn normal AI behavior and flag deviations immediately. Integrating training environments, interactive dashboards, and scalable diagnostic tools will enhance human ability to manage AI risks, ensuring better alignment with human values.
7. Limitations and Conclusion
One limitation of this work is the scope of the sampled and filtered papers. The rapidly growing literature on human-AI alignment spans diverse venues across many domains. Instead of an exhaustive collection, we focused on developing a holistic bidirectional human-AI alignment framework using essential research questions, dimensions, and codes. Our surveyed papers and team members primarily focus on computing-related fields like ML, NLP, and HCI, though alignment research also involves disciplines like cognitive science, psychology, and STS (Science, Technology, and Society). Our framework can naturally extend to these areas as needed. Despite these limitations, we believe our bidirectional human-AI alignment framework serves as a foundational reference for future researchers.
In conclusion, this study clarifies the definitions and scope of core terminologies of human-AI alignment and conducts a systematic review of over 400 related papers spanning diverse domains such as NLP, AI, HCI, and social science. Additionally, we introduce a novel conceptual framework of “Bidirectional Human-AI Alignment”, structuring the surveyed literature taxonomies into “aligning AI to humans” and “aligning humans to AI” with detailed categories and example papers. Furthermore, we identify limitations and risks in this area quantitatively and qualitatively, analyzing a fine-grained human value taxonomy, interaction modes for alignment, and discrepancies between AI and human evaluation. To pave the way for future studies, we discuss five stages to achieve the alignment goals from near-term to long-term perspectives and identify new possibilities to highlight future directions and opportunities in research.
APPENDIX
Appendix A Systematic Literature Review
A.1. Systematic Literature Review Process
To understand the research literature relevant to the ongoing, mutual process of human-AI alignment, we performed a systematic literature review based on the PRISMA guideline (Page et al., 2021; Stefanidi et al., 2023). Figure 7 shows the workflow of our process for paper coding and developing the bidirectional human-AI alignment framework. We introduce the step details below.
A.1.1. Identification and Screening with Keywords.
We started with papers published in the AI-related domain venues (including NLP, HCI, and ML fields) beginning from the advent of general-purpose generative AI to present, i.e., primarily between January, 2019 and January, 2024 (see details in Appendix A.2). We retrieved 34,213 papers in the initial Identification stage. Further, we collectively defined a list of keywords (see details in Appendix A.3) and screened for papers that included at least one of these keywords (e.g., human, alignment) or their variations either in the title or abstract. We included 2,136 papers in this Screening stage.
A.1.2. Assessing Eligibility with Criteria
We further filtered the 2,136 papers based on explicit inclusion and exclusion criteria, i.e., the Eligibility stage. Our criteria revolved around six research questions that we collectively identified to be most pertinent to the topic, including 1) what essential human values have been aligned by some AI models? 2) how did we effectively quantify or model human values to guide AI development? 3) what strategies have been employed to integrate human values into the AI development process? 4) how did existing studies improve human understanding and evaluation of AI alignment? 5) what are the practices for designing interfaces and interactions that facilitate human-AI collaboration? 6) How have AI been adapted to meet the needs of various human value groups? We included papers that could potentially answer any of these questions. Further, based on the scope in Section 3.4, we excluded papers that did not meet our inclusion criteria. This resulted in a final corpus of 411 papers, which were analyzed in detail using qualitative coding (see Appendix A.4 for more details).
A.1.3. Qualitative Code Development.
Referring to the code development process in Lee et al. (2024), we first conducted qualitative coding for each paper by identifying relevant sentences that could answer the above research questions, and entering short codes to describe them into a codebook. We iteratively coded relevant sentences from each paper through a mix of inductive and deductive approaches, which allowed flexibility to expand, modify or change the driving research questions based on our learnings as we went through the process. To ensure rigor in our coding process, two authors coded each paper. The first author independently annotated all papers after reviewing the paper abstracts and introductions. Twelve team members each annotated a subset of the paper corpus. Our corpus includes papers from different domains (e.g., HCI, NLP and ML). Therefore, we divided the authors into HCI and NLP/ML [^2] Note that NLP and ML are two different domains, we combine them together for the purposes of literature review analysis since they both work on developing and evaluating AI technologies. teams and assigned the papers accordingly based on expertise. All team members coded each of their assigned papers to answer all six questions (if applicable) introduced above.
A.1.4. Framework Development and Rigorous Coding.
After developing annotations, all authors collaborated to create the bidirectional human-AI alignment framework by integrating the annotations within each of the codes. The initial version of the framework was proposed by the author who reviewed all papers. This framework furthermore underwent iterative improvement through: 1) discussions with all team members involved in paper coding, and 2) revisions based on feedback from the project advisors. Additionally, we strengthened the framework by reviewing papers from the AI Ethics conferences (including FAccT and AIES), and related work of the collected papers that covered other domains such as psychology and social science. We further added missing codes and papers to ensure comprehensive coverage (see Appendix A.2 for details). The final bidirectional human-AI alignment framework, with detailed topologies, is presented in Section 4. Following the framework’s finalization, we conducted another separate coding process to annotate whether each paper investigated dimensions within our framework. Two authors independently coded each paper. [^3] The joint probability of agreement for the paper annotations was 0.78. These codes were then used to perform quantitative and qualitative analyses, as presented in Section 5.
Figure 7. The selection and refinement process of our systematic literature review. We referred to the PRISMA guideline (Page et al., 2021; Stefanidi et al., 2023) to report the workflow. From the identification of 34,213 records by keyword search, to screen eligible papers against our criteria and arriveg at our final corpus of 411 papers. For each of the stages where literature reviews were excluded (identification, screening, and eligibility) we further present the total of excluded records.
[…]
A.2. Venues
We primarily focused on papers from the fields of HCI, NLP, and ML ranging from year 2019 to 2024 January. We included all their papers tracks (e.g., CSCW Companion and Findings) without including workshops of conferences. From the ACL Anthology, OpenReview and ACM Digital Library, we retrieved 34,190 papers into a Reference Manager Tool (i.e., Paperpile). Particularly, the venues we surveyed are listed below.
  • HCI: CHI, CSCW, UIST, IUI;
  • NLP: ACL, EMNLP, NAACL, Findings
  • ML: ICLR, NeurIPS
  • Others: ArXiv, FAccT, AIES, and other related work
Additionally, we also consolidate the framework by reviewing the papers published in FAccT and AIES (i.e., important venues for AI Ethics research) between 2019 and 2024 and supplemented the codes, including the AI Regulatory and Policy code in Section 4.4.2 and the exemplary paper of Regulating ChatGPT (Hacker et al., 2023)), which were not covered by the original collections. Also, we include a number of papers in the “Other” class are found by related work that are highly relevant to this topic.
A.3. Keywords
We decided on a list of keywords relevant to bidirectional human-AI alignment. The detailed keywords include:
  • Human: Human, User, Agent, Cognition, Crowd
  • AI: AI, Agent, Machine Learning, Neural Network, Algorithm, Model, Deep Learning, NLP
  • LLM: Large Language Model, LLM, GPT, Generative, In-context Learning
  • Alignment: Align, Alignment
  • Value: Value, Principle
  • Trust: Trust, Trustworthy
  • Interact: Interact, Interaction, Interactive, Collaboration, Conversational
  • Visualize: Visualization, Visualize
  • Explain: Interpretability, Explain, Understand, Transparent
  • Evaluation: Evaluate, Evaluation, Audit
  • Feedback: Feedback
  • Ethics: Bias, Fairness
A.4. Inclusion and Exclusion Criteria
To further filter the most relevant papers among the keyword-filtered 2136 papers, we identified the six most important research questions we are interested in. We primarily selected the potential papers that can potentially address these six questions after reviewing their title and abstracts. The six topics of research questions in our filtering include:
  1. RQ.1
    [human value category] What essential human values have been aligned by some AI models?
  2. RQ.2
    [quantify human value] How did we effectively quantify or model human values to guide AI development?
  3. RQ.3
    [integrate human value into AI] What strategies have been employed to integrate human values into the AI development process?
  4. RQ.4
    [assess / explain AI regarding human values] How did existing studies improve human understanding and evaluation of AI alignment?
  5. RQ.5
    [human-AI interaction techniques] What are the practices for designing interfaces and interactions that facilitate human-AI collaboration?
  6. RQ.6
    [adapt AI for diverse human values] How has AI been adapted to meet the needs of various human value groups?
Particularly, we provide elaborated inclusion and exclusion criteria during our paper selection as listed below. We are aware that we have limitations during our paper filtering process.
Inclusion Criteria:
  • [[[[Human values]]]] we include papers that study human value definition, specification and evaluation in AI systems.
  • [[[[AI development techniques]]]] We include techniques of developing AI that aim to be more consistent with human values with interactions along all AI development stages (e.g., data collection, model construction, etc.)
  • [[[[AI evaluation, explanation and utilization]]]] we include papers that build human-AI interactive systems or conduct human studies to better evaluate, explain, and utilize AI systems.
  • [[[[building dataset with human interaction]]]] especially responsible dataset.
Exclusion Criteria:
  • [[[[Alignment not between human & AI]]]] we do not include alignment studies that are not between human and AI, such as entity alignment, cross-lingual alignment, cross-domain alignment, multi-modal alignment, token-environment alignment, etc.
  • [[[[AI models beyond LLMs - Modality]]]] we do not focus on AI models other than LLMs (e.g., 3D models, VR/AR, voice assistant, spoken assistant), our primary model modality is text. Specifically, we do not consider audio / video data; we do not consider pure computer vision modality.
  • [[[[No human-AI interaction]]]] we do not consider studies that do not involve the interaction between human and AI, such as (multi-agent) reinforcement learning. Specifically, we do not consider interactions via voices/speech, Do not consider game interaction; Do not consider interaction for Accessibility; Do not consider Mobile interaction; Not consider autonomous vehicle interaction wearable devices, or Physical interaction;
  • [[[[Tasks]]]] art and design, emotion.
  • [[[[No human included]]]]
  • [[[[focus on English]]]] primarily focus on English as the main language;
  • [[[[Application]]]] not include the NLP papers tailored for a specific traditional task, such as translation, entity recognition, sentiment analysis, knowledge graph, adversarial and defense, topic modeling, detecting AI generations, distillation, low resource, physical robots, text classification, games, image-based tasks, hate speech detection, Human Trafficking, etc.
  • [[[[Visualizing Embeddings]]]] Visualizing/interacting transformer embeddings?
  • [[[[Embedding-based]]]] explanation, evaluation, etc.
  • [[[[multi-agent reinforcement learning with self-play and population play]]]] techniques, such as self-play (SP) or population play (PP), produce agents that overfit to their training partners and do not generalize well to humans.
We acknowledge the extensive scope and rapid advancements of research in this area, and posit that our study offers insights that can be generalized to various modalities. For example, the value taxonomy and human-in-the-loop evaluation paradigm outlined in our framework can be applied to both text-based and other modality-based (e.g., vision, robotics) models. It’s worth noting that our literature review does not aim to exhaustively cover all papers in the field, which is impossible given the rapid advancement of human-AI alignment research. Instead, we adopt a human-centered perspective to review more than 400 key studies in this domain, focusing on delineating the framework landscape, identifying limitations, future directions, and a roadmap to pave the way for future research.
Appendix B Author contributions
This project was a team effort, built on countless contributions from everyone involved. To acknowledge individual authors’ contributions and enable future inquiries to be directed appropriately, we followed the ACM’s policy on authorship (of Publications, 2024) and listed contributors for each part of the paper below.
B.1. Overall Author List and Contributions
Project Lead
The project lead initialized and organized the project, coordinated with all authors, participated in the entire manuscript.
  • Hua Shen (University of Michigan, huashen@umich.edu): Initiated and led the overall project, prepared weekly project meetings, filtered papers, designed dimensions and codes (initial, revision), coded all papers, initiated the framework and developed human value and interaction modes analysis figures, participated in drafting all sections, paper revision and polishing.
Team Leads
The team leads organized all team events, coordinated with leads and members, contributed to a portion of manuscript.
  • Tiffany Knearem (Google, tknearem@google.com): Led the HCI team, prepared weekly team meetings, filtered papers, designed dimensions and codes (initial, revision), coded partial papers, ideated the framework and analysis and future work content, participated in writing (Critical Thinking and AI Impact on Human sections), paper revision and polishing.
  • Reshmi Ghosh (Microsoft, reshmighosh@microsoft.com): Led the NLP/AI team, prepared weekly team meetings, filtered papers, coded partial papers, ideated the framework and analysis and future work content, participated in writing (AI evaluation section), paper revision and polishing.
Team Members (Alphabetical)
The team members contributed to a portion of paper review, regular discussions, and drafted a portion of the manuscript.
  • Kenan Alkiek (University of Michigan, kalkiek@umich.edu): filtered papers, coded partial papers, data processing and analysis, ideated paper analysis and future work, paper revision and polishing, mainly involved in NLP Team
  • Kundan Krishna (Carnegie Mellon University, kundank@andrew.cmu.edu): filtered papers, coded partial papers, ideated the framework and future work, participated in writing (Customizing AI section), designed dimensions and codes (initial, revision), paper revision and polishing, mainly involved in NLP Team
  • Yachuan Liu (University of Michigan, yachuan@umich.edu): filtered papers, coded partial papers, participated in writing (revised Integrate General Value and Customization content sections), paper revision and polishing, mainly involved in NLP Team
  • Ziqiao Ma (University of Michigan, marstin@umich.edu): filtered papers, coded partial papers, designed dimensions and codes (initial, revision), developed Human Value category, participated in writing (Human Value taxonomy, revised representation, and value gap analysis sections), paper revision and polishing, mainly involved in NLP Team
  • Savvas Petridis (Google PAIR, petridis@google.com): filtered papers, coded partial papers, ideated the interaction-related analysis and future work, participated in writing (Perceive and Understand AI), paper revision and polishing, mainly involved in HCI Team
  • Yi-Hao Peng (Carnegie Mellon University, yihaop@cs.cmu.edu): filtered papers, coded partial papers, participated in writing (Human-AI Collaboration section), paper revision and polishing, mainly involved in HCI Team
  • Li Qiwei (University of Michigan, rrll@umich.edu): filtered papers, coded partial papers, ideated the interaction-related taxonomy and analysis, participated in writing (Interaction Mode section), mainly involved in HCI Team
  • Sushrita Rakshit (University of Michigan, sushrita@umich.edu): filtered papers, coded partial papers, participated in writing (Integrate General Value section), paper revision and polishing, mainly involved in NLP and HCI Team
  • Chenglei Si (Stanford University, clsi@stanford.edu): filtered papers, coded partial papers, designed dimensions and codes (initial, revision), ideated the framework and future work, participated in writing (Assessment of Collaboration and Impact section), paper revision and polishing, mainly involved in HCI Team
  • Yutong Xie (University of Michigan, yutxie@umich.edu): filtered papers, coded partial papers, designed dimensions and codes (initial, revision), ideated the value representation taxonomy, participated in writing (Human Value Representation section), paper revision and polishing, , mainly involved in NLP Team
Advisors (Alphabetical)
The advisors involved in and made intellectual contributions to essential components of the project and manuscript.
  • Jeffrey P. Bigham (Carnegie Mellon University, jbigham@cs.cmu.edu): contributed to the framework on aligning human to AI direction, vision on the status quo of alignment research, and future work discussions, and participated in paper revision and proofreading.
  • Frank Bentley (Google, fbentley@google.com): contributed to the historical context and project objectives, improved the definitions and design of research methodology, and participated in paper revision and proofreading.
  • Joyce Chai (University of Michigan, chaijy@umich.edu): iteratively involved in developing and revising definitions and the framework on aligning AI to human direction, advised on analysis and future work, and participated in paper revision and proofreading.
  • Zachary Lipton (Carnegie Mellon University, zlipton@cmu.edu): contributed insights from Machine Learning, NLP, and AI fields to revise the definitions and framework on aligning AI to human direction, and participated in paper revision and proofreading.
  • Qiaozhu Mei (University of Michigan, qmei@umich.edu): contributed insights from Data Science, Machine Learning, and NLP fields to improve definitions and the framework on aligning AI to human direction, and participated in paper revision and proofreading.
  • Rada Mihalcea (University of Michigan, mihalcea@umich.edu): involved in framing and revising the structure and taxonomy of human values, and contributed to improving the manuscript’s title, introduction, and other sections, and participated in paper revision and proofreading.
  • Michael Terry (Google Research, michaelterry@google.com): contributed arguments and vision on the status quo of alignment research, framed project objectives and contributions, improved definitions and data analysis, and participated in paper revision and proofreading.
  • Diyi Yang (Stanford University, diyiy@stanford.edu): involved in improving definitions and the framework, contributed social insights to the work, and participated in paper revision and proofreading.
Project Leading Advisors
The project leading advisors actively involved in the entire project process and all manuscript sections.
  • Meredith Ringel Morris (Google DeepMind, merrie@google.com): iteratively involved in drafting all sections, contributed to core argument ideation, framework and definition improvement, provided future work insights, and participated in paper drafting, revision, and proofreading on all sections.
  • Paul Resnick (University of Michigan, presnick@umich.edu): actively involved and advised on the entire project process, including initiating the project and research agenda, iteratively improved definitions, framework, and analysis, and participated in paper revision and proofreading.
  • David Jurgens (University of Michigan, jurgens@umich.edu): provided advice throughout the project, including iterative discussions on project milestones and content ideation, organized several meetings to receive feedback from external audiences, and participated in paper revision and proofreading.
Acknowledgements.
We thank Eric Gilbert for his constructive feedback on human-centered insights on alignment, thank Eytan Adar for his valuable guidance on designing the interaction techniques for human-AI alignment, and thank Elizabeth F. Churchill for her insightful discussion on this manuscript. We also thank Michael S Bernstein, Denny Zhou, Cliff Lampe, and Nicole Ellison for their encouraging feedback on this work. We welcome researchers’ constructive discussions and interdisciplinary efforts to achieve long-term and dynamic human-AI alignment collaboratively in the future. This work was supported in part by the National Science Foundation under Grant No. IIS-2143529 and No. IIS-1949634.

### v4

@@
-Towards Bidirectional Human-AI Alignment: A Systematic Review for Clarifications, Framework, and Future Directions
-Abstract.
-Recent advancements in general-purpose AI have highlighted the importance of guiding AI systems towards the intended goals, ethical principles, and values of individuals and groups, a concept broadly recognized as alignment.
-However, the lack of clarified definitions and scopes of human-AI alignment poses a significant obstacle, hampering collaborative efforts across research domains to achieve this alignment.
-In particular, ML- and philosophy-oriented alignment research often views AI alignment as a static, unidirectional process (i.e., aiming to ensure that AI systems’ objectives match humans) rather than an ongoing, mutual alignment problem (Wikipedia, 2024).
-This perspective largely neglects the long-term interaction and dynamic changes of alignment.
-To understand these gaps, we introduce a systematic review of over 400 papers published between 2019 and January 2024, spanning multiple domains such as Human-Computer Interaction (HCI), Natural Language Processing (NLP), Machine Learning (ML).
-We characterize, define and scope human-AI alignment.
-From this, we present a conceptual framework of “Bidirectional Human-AI Alignment” to organize the literature from a human-centered perspective.
-This framework encompasses both 1) conventional studies of aligning AI to humans that ensures AI produces the intended outcomes determined by humans, and 2) a proposed concept of aligning humans to AI, which aims to help individuals and society adjust to AI advancements both cognitively and behaviorally.
-Additionally, we articulate the key findings derived from literature analysis, including literature gaps and trends, human values, and interaction techniques.
-To pave the way for future studies, we envision three key challenges and give recommendations for future research.
-Human-AI Alignment, Human-AI Interaction
-copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Manuscript; submitted to ACM; 2024ccs: Human-centered computing HCI theory, concepts and models
-1. Introduction
-Artificial Intelligence (AI) has advanced significantly, especially with the advent of general-purpose generative AI, demonstrating unprecedented capabilities in solving a wide range of complicated and challenging problems, such as reasoning, generation, language understanding, and more (Ouyang et al., 2022).
-However, as AI becomes increasingly powerful and integrated into daily life, society is confronted with an array of risks to both individuals and to broader society (Morris et al., 2024).
-For example, text-to-image generative models were found to amplify stereotypes about race and gender (blo, 2023), and biased algorithms in hiring processes were found to perpetuate discrimination (Park et al., 2021a).
-These risks illuminate foundational questions around how and which values are embedded in AI models that drive decisions within real-world contexts.
-[…]
-[…]
-[…]
-Figure 1.
-The overview of the Bidirectional Human-AI Alignment framework.
-Our framework encompasses both Refer to caption conventional studies of “Align AI to Humans” in AI development that ensures AI produces the intended outcomes determined by humans, and  Refer to caption a novel concept of “Align Humans to AI” , which aims to help humans and society to better understand, critique, collaborate, and adapt to transformative AI advancements.
-Note that “Humans” refers to both AI users and those who do not interact with AI but may be impacted by AI systems.
-We further identify four key research questions (i.e., RQ1-RQ4) to facilitate this holistic loop of “bidirectional human-AI alignment”, and organize the literature that can potentially address RQ1-RQ4 in Section 4.
-This paper conducts a systematic literature review of over 400 human-AI alignment papers across multiple disciplines, based on PRISMA guidelines (Page et al., 2021; Stefanidi et al., 2023).
-The literature review corpus includes papers drawn from the the Human-Computer Interaction (HCI), Natural Language Processing (NLP) and Machine Learning (ML) domains that were published between the advent of general-purpose generative AI to present, i.e., primarily between January 2019 and January 2024.
-We take an interdisciplinary approach to human-AI alignment (Gilbert, 2024), drawing from theories and empirical studies across the aforementioned domains.
-To round out our approach, we composed our research team of members from the domains of HCI, NLP, ML, Data Science, Computational Social Science, and Cognitive Science.
-    Definition.
-Bidirectional Human-AI Alignment is a comprehensive framework that encompasses two interconnected alignment processes: ‘Aligning AI to Humans’ and ‘Aligning Humans to AI’.
-The former focuses on integrating human specifications to train, steer, and customize AI, while the latter investigates human cognitive and behavioral adaptations to AI, which supports humans in understanding, critiquing, collaborating with, and adapting to AI advancements.
-Paper Organization and Key Findings.
-We provide clarified definitions and scopes related to human-AI alignment, including “what is the goal of alignment?”, “with whom to align?”, and “what values should be aligned with?”(Section 3).
-We then present a conceptual framework of Bidirectional Human-AI Alignment from a long-term and dynamic perspective, encompassing both “Align AI to Humans” and “Align Humans to AI” (see Figure 1).
-Furthermore, we identify four key research questions (RQs) in the Bidirectional Human-AI Alignment framework and provide a structured way to organize the existing research literature to address these questions in Section 4.
-The resulting structured topologies in Figure 2 and 3 aim to provide a shared vocabulary that can help streamline communication and collaboration between alignment researchers in different disciplines.
-Furthermore, through iterative paper coding and literature analysis, we derive insights including meta analysis on gaps and trends in existing alignment research, findings of human values, and the interaction techniques for human-AI alignment (Section 5).
-To pave the way for future studies, we further envision three future directions in Figure 6 and Section 6.
-Our study highlights several key findings.
-For instance, many essential human values outlined in Table 2, such as loyalty and environmental protection, are crucial for alignment but are often overlooked in current research.
-Additionally, many topics within the bidirectional alignment framework, including leveraging implicit human feedback to align AI models and auditing AI for various social values, are important but underexplored.
-Operationalizing the Framework.
-To help AI practitioners (e.g., developers, researchers, user experience designers) who are interested in this field to leverage the bidirectional approaches, we highlight some systematic resources about human-AI alignment provided both within this survey and on companion Github repository:
-  • A Github repository of “Bidirectional Human-AI Alignment” maintaining latest reading lists and relevant updates.
-  • Fig. 1: An overview of the bidirectional human-AI alignment framework with formal definition.
-  • Table 1: Typical alignment goals, their definitions, and limitations.
-  • Table 2: A comprehensive list of basic human values, their relationships, and related alignment papers.
-  • Fig. 2 and 3: Fine-grained topologies and papers of “Aligning AI to Humans” & “Aligning Humans to AI” research.
-  • Fig. 4: An illustration of trends and gaps of alignment research dimension distribution.
-  • Fig. 5: A summary of common interactive techniques to specify human values and exemplary papers.
-2. Challenges in Achieving Alignment
-[…]
-Challenge 1: Outer and Inner Alignment.
-In the context of “intelligent agents,” until now, AI alignment research has aimed to ensure that any AI systems that would be set free to make decisions on our behalf would act appropriately and reduce unintended consequences (Wooldridge, 1999; Stuurman and Wijnands, 2001; Russell and Norvig, 2016).
-At the near-term stage, aligning AI involves two main challenges: carefully specifying the purpose of the system (outer alignment i.e., providing well-specified rewards (Ngo et al., 2024)) and ensuring that the system adopts the specification robustly (inner alignment, i.e., ensuring that every action given an agent in a particular state learns desirable internally-represented goals (Ngo et al., 2024)).
-Significant efforts have been made, for inner alignment, to align AI systems to follow alignment goals of an individual or a group (e.g., instructions, preferences, values, and/or ethical principles) (Ouyang et al., 2022) and to evaluate the performance of alignment (Santurkar et al., 2023).
-However, for outer alignment, AI designers are still facing difficulties in specifying the full range of desired and undesired alignment goals of humans.
-Challenge 2: Specification Gaming.
-To learn human alignment goals, AI designers typically provide an objective function, instructions, reward function, or feedback to the system, which is often unable to completely specify all important values and constraints that a human intended (Hemphill, 2020).
-Hence, AI designers resort to easy-to-specify proxy goals such as maximizing the approval of human overseers (Wikipedia, 2024), which results in “specification gaming” (Krakovna et al., 2020) or “reward hacking” (Pan et al., 2022) issues (i.e., AI systems can find loopholes that help them accomplish the specific objective efficiently but in unintended, possibly harmful ways).
-Additionally, the black-box nature of neural networks brings additional ethical and safety concerns for alignment because humans don’t know about the inner state and the actions AI leveraged to achieve the output.
-Consequently, AI systems might make “correct” decisions with “incorrect” reasons, which are difficult to discern.
-Society is already facing these issues, such as data privacy (Tucker et al., 2018), algorithmic bias (Hajian et al., 2016), self-driving car accidents (Bonnefon et al., 2016), and more.
-As a result, these considerations necessitate considering human-AI interaction in AI alignment for specification and evaluation, ranging from addressing problems around who uses an AI system, with what goals to specify, and if the AI system perform its intended function from the user’s perspective (Weidinger et al., 2023).
-Challenge 3: Scalable Oversight.
-From a long-term perspective, when advanced AI systems become more complex and capable (e.g., AGI (Morris et al., 2024)), it becomes increasingly difficult to align them to human values through human feedback.
-Evaluating complex AI behaviors applied to increasingly challenging tasks can be slow or infeasible for humans to ensure all sub-steps are aligned with their values (Terry et al., 2023).
-Therefore, researchers have begun to investigate how to reduce the time and effort for human supervision, and how to assist human supervisors, referred to as Scalable Oversight (Amodei et al., 2016).
-Challenge 4: Dynamic Nature.
-As AI systems become increasingly powerful, the alignment solutions must also adapt dynamically since human values and preferences change as well.
-As Dautenhahn et al.
-(2000) posit, AI systems may be neither humane nor desirable if we do not ask questions about the long-term cognitive and social effects of social agent systems (e.g., how will agent technology affect human cognition).
-All these considerations call for a long-term and dynamic perspective to address human-AI alignment as an ongoing, mutual process with the collective efforts of cross-domain expertise.
-Challenge 5: Existential Risk.
-Further, some AI researchers claim that (Bengio et al., 2024) advanced AI systems will begin to seek power over their environment (e.g., humans) once deployed in real-world settings, as such behavior may not be noticed during training.
-For example, some language models seek power in text-based social environments by gaining money, resources, or social influence (Pan et al., 2023).
-Consequently, some hypothesize that future AI, if not properly aligned with human values, could pose an existential risk to humans (Dafoe and Russell, 2016).
-Table 1.
-The Goals of Alignment.
-We present the six prevailing alignment goals, associating with their Definitions (middle column), Limitations and Risks (right column).
-We consider Human Values as the main goal of alignment in this work referring to an extensive analysis and arguments in existing studies (Gabriel, 2020; Russell, 2014)
-3. Fundamental Definitions and Clarifications
-Addressing alignment challenges is a complex and multifaceted process.
-To develop a comprehensive perspective of the ongoing and mutual process of human-AI alignment, this section begins by introducing core definitions and key components of alignment.
-Subsequently, we outline our methodology for conducting a systematic literature review and the rationale behind our proposed framework of bidirectional human-AI alignment.
-3.1. What is the goal of alignment?
-[…]
-Table 2.
-The relations and a fine-grained taxonomy of 69 exemplary human values.
-We consider 5 high-order value types encompassing 12 motivational value types, indicated by their sources (e.g., individuals, society and interaction).
-The exemplary values with red dot ([Uncaptioned image]) indicates there are no work in our surveyed papers examining the specific values.
-3.2. With whom to align?
-Pertinent stakeholders within the AI landscape can be the potential objects for AI to align with, including lay people (e.g., end users of AI systems) (Qi et al., 2023), AI practitioners (e.g., developers, researchers) (Boggust et al., 2022; Anwar et al., 2024; Ngo et al., 2024), organizational entities (e.g., technology firms, professional communities), national/international bodies (e.g., governments, legislative bodies) (Deshpande and Sharp, 2022) and others.
-Many alignment research papers have focused on general humans without specifying particular groups (Yao et al., 2023b; Liu et al., 2022; Guo et al., 2024).
-Nevertheless, different groups hold different, sometimes even contrasting, values (Gabriel, 2020).
-As a consequence, rather than identifying a true moral theory as a one-size-fits-all value, prior alignment research argues to select the appropriate principles for compatible human groups (Sorensen et al., 2024).
-To this end, pluralistic value alignment, grounded on social choice theory (Arrow, 2012), proposes combining individual views fairly in developing alignment principles as a potential solution (Sorensen et al., 2024).
-In this work, we also consider values from this pluralistic perspective, where AI should be aligned with pluralistic human individuals and societal groups who would ultimately be impacted by AI.
-3.3. What are the values to be aligned with?
-[…]
-A Comprehensive Taxonomy of Human Values.
-Nevertheless, this conventional theory was developed without the context of human-AI interaction, which might overlook values that need to be considered for human-AI alignment.
-Therefore, we used a bottom-up approach to extract all values studied in our collected alignment literature (elaborated in Section A.1), mapped them onto the Schwartz Theory of Basic Values, and supplemented the theory with AI-related structure and content.
-As a result, we identified the structural relationships among human values (see Figure LABEL:fig:value_structure) and mapped existing literature to a fine-grained taxonomy (see Table 2).
-As shown in Figure LABEL:fig:value_structure, we supplemented the traditional theory’s four high-order value types (i.e., “Self-Enhancement”, “Openness to Change”, “Conservation”, “Self-Transcendence”) with a novel high-order value type, named “Desired Values for AI Tools” that encompasses two motivational value types (i.e., “Usability” and “Human-Likeness”).
-We further organize the relationship among these value types along two dimensions (Schwartz, 2012): different resources (i.e., individuals, society and interaction) and different self-intentions (i.e., self-protection against threat and self-expansion and growth).
-Furthermore, we elaborate the definitions of the 12 motivational value types and their exemplary values by mapping them to relevant human-AI alignment papers from our corpus in Table 2.
-During the process of mapping, we found: 1) value terms in empirical papers were often named differently (e.g., capability and competence), or check their opposites (e.g., fairness and bias); 2) there are many values not studied in our corpus, i.e., indicated as ([Uncaptioned image]) in the Figure.
-    Takeaways & Implications for Alignment Clarifications 1.
-We propose the “pluralistic human values” as the alignment goal, wherein AI should be aligned with the diverse perspectives of individuals and societal groups who may be directly or indirectly impacted by AI.
-2.
-Many exemplary human values outlined in Table 2, such as loyalty and environmental protection, are important for alignment but are often overlooked or underexplored in current research.
-3.
-The human value system should account for complex application contexts and relationships, such as priorities and conflicts, rather than focusing on an exclusive subset of values.
-3.4. Scopes and Key Components in Alignment
-This section clarifies the key components we considered in human-AI alignment. Additionally, we define the scopes of each key component as commonly adopted in current research.
-  • Humans.
-Our primary focus is on human individuals, groups, or organizations that will ultimately develop, use, and potentially impact or be influenced by AI systems, as these are the entities with which AI systems should align.
-We emphasize the importance of considering the pluralistic values of diverse users rather than treating users as a monolithic group.
-  • Artificial Intelligence (AI).
-We focus on AI systems including both domain-specific AI systems that address specific tasks (e.g., reasoning, dialogue) and general-purpose AI models that aim to complete any tasks with performance comparable to a human’s.
-These AI systems include generative, classification, and regression models, among others.
-Particularly, we primarily focus on language models as the representative AI models for alignment research, and discuss how the insights from this study can be generalized to other modalities at the section end.
-  • Alignment.
-Our review encompassed all the alignment goals outlined in Section 3.1.
-Since many studies emphasized the importance of value alignment (Gabriel, 2020; Russell, 2014), we particularly summarized a clarified taxonomy of alignment values and identified the value in each paper (if applicable).
-Additionally, we focus on analyzing the AI models’ output and generation, but not the neural network’s intermediate representations (McGrath et al., 2023; Park et al., 2023; Xie et al., 2024), for alignment research.
-3.5. Systematic Literature Review and Framework Rationale
-Achieving the alignment between AI and human values necessitates collaboration among interdisciplinary AI practitioners (e.g., developers and researchers).
-To comprehensively analyze the research literature pertaining to this ongoing, mutual process of human-AI alignment, we conducted a systematic literature review adhering to the the PRISMA guideline (Page et al., 2021; Stefanidi et al., 2023).
-Our iterative selection process yielded 411 papers from an initial pool of 34,213 publications, spanning the period from January 2019 to January 2024.
-These papers were sourced from leading venues across multiple AI-related domains, including Natural Language Processing, Human-Computer Interaction, Machine Learning, and more, using well-defined keywords and criteria.
-A group of 13 interdisciplinary researchers coded the papers.
-Through iterative discussions, we developed the paper codes and proposed framework (Section 3).
-Detailed steps and processes are elucidated in Appendix A.
-3.6. Rationale Behind the Framework
-Our systematic review revealed that current research largely overlooks the nature of long-term dynamic changes and the importance of interaction in human-AI alignment.
-First, the manifestation of certain AI risks, harms, and the practical quality of AI services primarily emerge through interaction between humans and AI systems (Weidinger et al., 2023).
-For instance, Shen and Huang (2020) discovered that AI explanations, deemed effective by AI developers and researchers, were found not useful for humans in real-world scenarios of interactive AI error detection.
-Additionally, the use of AI influences human thinking and behavior, which in turn alters their values when interacting with AI.
-For instance, Li et al.
-(2024b) found that incorporating AI assistance in writing workflows changed humans’ writing perceptions, behavior, and performance, creating new requirements for developing the next generation of AI writing assistants (Lee et al., 2024).
-In light of these findings, we posit that AI practitioners must expand their focus beyond merely “Aligning AI to Humans” – a paradigm centered on integrating human specifications to train, steer, and customize AI.
-Equal emphasis must be placed on “Aligning Humans to AI”, a complementary paradigm investigating human cognitive and behavioral adaptation to AI.
-This latter aspect is crucial in supporting individuals and society at large in understanding, critiquing, collaborating with, and adapting to transformative AI advancements.
-To encapsulate these comprehensive perspectives, we propose the “Bidirectional Human-AI Alignment” framework, and provide its definition below Figure 1.
-This holistic framework acknowledges the reciprocal nature of alignment, addressing both the advancements of AI systems to incorporate human values, and the evolving human adaptation of AI systems in real-world practice.
-4. Bidirectional Human-AI Alignment Framework
-Figure 2.
-The fine-grained topology of “Align AI to Humans” direction, which studies mechanisms to ensure that AI system’s objectives match those of humans’.
-The goal is to integrate human specification to train, steer and customize AI systems.
-This section introduces the Bidirectional Human-AI Alignment framework, which encompassing two interconnected alignment directions, as shown in Figure 1.
-The [Uncaptioned image] “Align AI to Humans” direction studies mechanisms to ensure that AI systems’ values match those of humans’ (Section 4.1 and 4.2).
-We identify two crucial research questions to be addressed in this direction: RQ1.
+Position: Towards Bidirectional Human-AI Alignment
+Abstract
+Recent advances in general-purpose AI underscore the urgent need to align AI systems with human goals and values.
+Yet, the lack of a clear, shared understanding of what constitutes "alignment" limits meaningful progress and cross-disciplinary collaboration.
+In this position paper, we argue that the research community should explicitly define and critically reflect on "alignment" to account for the bidirectional and dynamic relationship between humans and AI.
+Through a systematic review of over 400 papers spanning HCI, NLP, ML, and more, we examine how alignment is currently defined and operationalized.
+Building on this analysis, we introduce the Bidirectional Human-AI Alignment framework, which not only incorporates traditional efforts to align AI with human values but also introduces the critical, underexplored dimension of aligning humans with AI – supporting cognitive, behavioral, and societal adaptation to rapidly advancing AI technologies.
+Our findings reveal significant gaps in current literature, especially in long-term interaction design, human value modeling, and mutual understanding.
+We conclude with three central challenges and actionable recommendations to guide future research toward more nuanced, reciprocal, and human-AI alignment approaches.
+1 Introduction
+Artificial Intelligence (AI), particularly generative AI, has demonstrated remarkable capabilities in reasoning, language understanding, problem solving, and more ouyang2022training .
+However, its increasing integration into society raises significant risks, such as amplifying biases in hiring bloomberg or perpetuating stereotypes in text-to-image models park2021human .
+These concerns highlight the urgent need to align these systems with values, ethical principles, and the goals of individuals and society at large.
+This need, commonly referred to as “AI alignment,” wiki_AI_alignment ; terry2023ai is crucial for ensuring that AI systems function in a manner that is not only effective but also consistent with human values, minimizing harm and maximizing societal benefits.
+Yet, key challenges remain:
+Challenge 1: Specification Gaming.
+AI designers often define objectives or feedback to align systems with human goals, but these rarely capture all intended values hemphill2020human .
+This leads to reliance on proxies like human approval wiki_AI_alignment , enabling specification gaming krakovna2020specification ; pan2022effects , where AI makes seemingly “correct” decisions for the wrong, opaque reasons tucker2018privacy ; hajian2016algorithmic ; bonnefon2016social .
+Challenge 2: Scalable Oversight.
+As AI systems become more complex—potentially reaching AGI morris2024levels —aligning them through human feedback grows harder.
+Evaluating their behavior is often slow or infeasible terry2023ai , prompting research into reducing supervision burdens and enhancing human oversight, a challenge known as Scalable Oversight amodei2016concrete .
+Challenge 3: Dynamic Nature.
+As AI advances, alignment must adapt to evolving human values.
+Without considering long-term cognitive and social impacts, AI may become neither humane nor desirable dautenhahn2000living .
+This needs a dynamic, ongoing alignment process with cross-disciplinary collaboration.
+Figure 1: The overview of the Bidirectional Human-AI Alignment framework.
+Our framework encompasses both Refer to caption conventional studies which focus on “Align AI with Humans” and  Refer to caption “Align Humans with AI”.
+We further identify four key Research Questions to facilitate this holistic loop of “bidirectional human-AI alignment”, and provide answers that can potentially address RQ1-RQ4 in Section 3.
+Traditionally, AI alignment has been approached as a static, one-way process focused on shaping AI to achieve desired outcomes and avoid harm goyal2024designing ; ouyang2022training ; santurkar2023whose .
+However, this unidirectional view is increasingly insufficient as AI systems become more integrated into daily life and assume complex decision-making roles carroll2024ai .
+Their interactions with humans create evolving feedback loops that influence both AI behavior and human responses shen2020useful ; li2024value ; lee2024design , highlighting the need for a more dynamic and reciprocal understanding of alignment carroll2024ai .
+In this position paper, we argue that it is critical for the research community to explicitly reflect on what we mean by “alignment” and to take into account the bidirectional, dynamic interactions between humans and AI to achieve responsible and safe AI systems.
+Through a systematic review of over 400 papers across HCI, NLP, ML, and related fields, we examine how alignment is currently defined and implemented.
+Based on this analysis, we propose the Bidirectional Human-AI Alignment (in Figure 1) framework.
+It extends the traditional focus on “Aligning AI to Humans”—integrating human input to train, steer, and customize AI—by introducing the equally vital yet underexplored direction of “Aligning Humans to AI”, which emphasizes cognitive, behavioral, and societal adaptation to rapid AI advancement.
+Our findings reveal key gaps in existing research, particularly in human value modeling, oversight of model inference, critical evaluation of AI’s embedded values, and its broader societal impact.
+We conclude by outlining near- and long-term risks and opportunities, offering actionable recommendations to advance more reciprocal, adaptive, and nuanced approaches to human-AI alignment.
+2 Defining Alignment: Fundamentals
+Building on our analysis of systematic review (see details in Appendix 8), we explicitly identify the key definitions in alignment research and formally propose “Bidirectional Human-AI Alignment”.
+Goals.
+AI alignment research proposes multiple alignment goals park2023generative ; xie2024can , such as intentions ouyang2022training ; anwar2024foundational , preferences bansal2023peering ; verma2023hindsight , instructions bai2022constitutional ; liu2023makes , and values sorensen2024roadmap ; gabriel2020artificial .
+But these terms are often used interchangeably without clear distinctions.
+Philosophical analysis suggests human values (moral beliefs/principles) are the most suitable alignment goal, as they ensure AI acts ethically while minimizing risks gabriel2020artificial ; stuart2014whitepaper .
+Though trade-offs exist, this work adopts "human values" as the alignment objective, meaning AI should behave as individuals or society morally expect (See details in Table 2).
+Align with Whom.
+AI alignment involves multiple stakeholders, such as end users qi2023fine , AI practitioners boggust2022shared ; anwar2024foundational ; ngo2022alignment , and organizations deshpande2022responsible .
+Many studies reference “general humans” without accounting for group differences, despite the fact that stakeholders often hold conflicting values gabriel2020artificial .
+Rather than pursuing a universal moral theory, prior work advocates aligning AI with principles tailored to compatible human groups sorensen2024roadmap .
+Pluralistic value alignment, grounded in social choice theory arrow2012social ; sorensen2024roadmap , provides a framework for integrating diverse perspectives sorensen2024roadmap .
+We adopt this view, recognizing that aligning with affected individuals and groups entails ongoing challenges.
+Align with What.
+While prior studies have sought to align AI with human values, these values are often vaguely or inconsistently defined.
+To address this, we reviewed several prominent value theories, including Moral Foundations Theory graham2013moral and Social Norms and Ethics forbes2020social , drawing from psychology and social science.
+We selected the Schwartz Theory of Basic Values schwartz1994there ; schwartz2012overview for its demonstrated cross-cultural applicability, relevance across contexts (individuals, interactions, groups), and frequent adoption in NLP research kang2023values ; kiesel2022identifying .
+This framework consists of eleven types of universal values, and defines values as beliefs about desirable end states or behaviors that transcend specific situations and guide evaluation and decision-making.
+Definition: Bidirectional Human-AI Alignment is a comprehensive framework that encompasses two interconnected alignment processes: ‘Aligning AI with Humans’ and ‘Aligning Humans with AI’.
+The former focuses on integrating human specifications into training, steering, and customizing AI.
+The latter supports human agency, empowering people to think critically when using AI, collaborate effectively with it, and adapt societal approaches to maximize its benefits for humanity.
+3 Bidirectional Human-AI Alignment Framework
+Table 1: The fine-grained typology of two directions in the Bidirectional Human-AI Alignment.
+Research Question Sub-Research Question Dimensions References
+Align AI with Humans RQ1: Human Values and Specifications Categorizing Aligned Human Values what values have been aligned with AI?
+Source of Values kiesel2022identifying ; jin2022make ; ammanabrolu2022aligning
+Value Types santurkar2023whose ; nie2024moca ; santy2023nlpositionality
+Interaction Techniques to Specify AI Values How humans could interactively specify values in AI development? Explicit Human Feedback petridis2024constitutionmaker ; dai2023safe ; bai2022constitutional
+Implicit Human Feedback isajanyan2024social ; park2020scalable ; lu2022towards
+Simulated Human Value Feedback petridis2024constitutionalexperts ; dong2023aligndiff ; kim2023aligning
+RQ2: Integrating Human Specification into AI Develop AI with General Values how to incorporate general human values into AI development?
+Instruction Data shen2023multiturncleanup ; sun2024principle ; dubois2024alpacafarm
+Model Learning ouyang2022training ; dong2023raft ; rafailov2024direct
+Inference Stage bai2022constitutional ; gou2023critic ; bradley2023quality
+Customizing AI for Individuals or Groups how to customize AI to incorporate values of individuals or groups? Customized Data felkner2023winoqueer ; orlikowski2023ecological ; lyu2023exploiting
+Adapt Model by Learning zhao2023group ; peschl2022moral ; sun2022bertscore
+Interactive Alignment maghakian2022personalized ; swazinna2022user ; lahoti2023improving
+Evaluating AI Systems how to evaluate AI regarding human values? Human-in-the-loop Evaluation sharma2023cognitive ; ouyang2023shifted ; ammanabrolu2022aligning
+Automatic Evaluation jia2024towards ; sun2024principle ; patel2021stated
+Ecosystem how to build the ecosystem to facilitate human-AI alignment? Platforms dubois2024alpacafarm ; xu2023gentopia ; zhou2023sotopia ; yuan2024uni ; gerstgrasser2021crowdplay ; pei2022potato
+Align Humans with AI RQ3: Human Cognitive Adjustment to AI Perceiving and Understanding of AI how do humans learn to perceive and explain AI systems?
+Education and Training Human long2020ai ; ketan-ai-medicine-education ; mcdonald2020intersectional
+AI Sensemaking and Explanations kaur2022sensible ; sivaraman2023ignore ; tenney2020language
+Critical Thinking about AI how do humans think critically about AI systems? Trust and Reliance on AI Decisions yin2019understanding ; ma2023should ; schemmer2023appropriate
+Ethical Concerns and AI Auditing holstein2019improving ; bandy2021problematic
+Calibrate Cognition to Align AI he2023interaction ; kocielnik2019will ; wischnewski2023measuring
+RQ4: Human Adaptation Behavior to AI Human Collaborating with Diverse AI Roles how do humans collaborate with AI with differing capability levels?
+AI Assistants wu2022ai ; gebreegziabher2023patat ; scattershot
+AI Partners park2023generative ; zagalsky2021design ; bernstein2010soylent
+AI Tutors ma2023hypocompass ; peng2021say ; shaikh2023rehearsal
+AI Impacts on Humans and Society how are humans influenced by AI systems? Impact on Individual Behavior shen2020useful ; ashkinaze2024ai
+Societal Concerns and AI Impacts 10.1145/3610082 ; kazemitabaar2023novices
+Reaction to AI Advancements lima2023blaming ; hacker2023regulating
+Evaluation in Human Studies how might we evaluate the impact of AI on humans and society? Evaluate Human-AI Collaboration Lee2022EvaluatingHM ; ding-etal-2023-harnessing
+Evaluate Societal Impact santurkar2023whose ; Shin2023SuperhumanAI
+This section introduces the Bidirectional Human-AI Alignment framework, which encompasses two interconnected alignment directions as a feedback loop, as shown in Figure 1.
+The “Align AI to Humans” direction refers to mechanisms to ensure that AI systems’ values match those of humans’.
+The “Align Humans to AI” direction investigates human cognitive and behavioral adaptation to AI advancement.
+We introduce more details below.
+3.1 Align AI to Humans
+This direction delineates alignment research from AI-centered perspective (e.g., ML/NLP domains) and provides AI developers and researchers with approaches for handling two main challenges: carefully specifying the values of the system, and ensuring that the system adopts the specification robustly wiki_AI_alignment ; ngo2022alignment .
+Therefore, as shown in Table 1, we explore two core research questions in this direction as: RQ1.
 What relevant human values are studied for AI alignment, and how do humans specify these values?
 and RQ2.
-How can human values be integrated into the development of AI?
-In comparison, the [Uncaptioned image] “Align Humans to AI” direction investigates the humans’ cognitive and behavioral adaptation to the AI advancement (Section 4.3 and 4.4).
-We recognize two imperative research questions along this direction: RQ3.
-How might humans learn to perceive, explain, and critique AI?
-and RQ4.
-How do humans and society make behavioral changes and react to AI advancement?.
-We address the four questions by elucidating the fine-grained typologies and literature solutions derived from our systematic review, as delineated in Figure 2 and Figure 3, respectively.
-DIRECTION-I:  [Uncaptioned image] ALIGN AI to HUMANS
-This direction delineates alignment research from AI-centered perspective (e.g., ML/NLP domains) and provides AI developers and researchers with approaches for two main challenges: carefully specifying the values of the system, and ensuring that system adopts the specification robustly (Wikipedia, 2024; Ngo et al., 2024).
-Therefore, as shown in Figure 2, we explore the two core research questions in this direction as: RQ1.
-Human Values and Specifications (Section 4.1) and RQ2.
-Integrating Human Specifications into AI (Section 4.2)
-4.1. Align AI to Humans: [Uncaptioned image] Human Values and Specifications
-RQ1: What relevant human values are studied for AI alignment, and how do humans specify these values?
-We structure existing literature to address this research question by answering the following “Sub-Research Questions”: What values have been aligned by AI?
-(Section 4.1.1) and then exploring How humans could interactively specify values in AI development?
-(Section 4.1.2).
-As shown at the top of Figure 2, particularly, we articulate the “Dimensions“ we summarized to answer each of these sub-research questions, and provide “Codes“ associated with “Example Papers” that have been studied in each dimension.
-4.1.1. Categorizing Aligned Human Values. What values have been aligned with AI?
-To clarify and systematically understand human values relevant to human-AI alignment, we leverage the adapted “Schwartz Theory of Basic Values” introduced in Section 3.3, and examine the category of aligned human values from the two dimensions of “Sources” and “Types”.
-Sources of Values.
-This dimension examines the three sources of human values (Schwartz, 1994).
-Individual sources indicate values from individuals comprise universal needs of individuals as biological organisms (Schwartz, 1994) or prioritize personal interests (Kiesel et al., 2022; Jin et al., 2022).
-At this level, value alignment can be usually assessed independently of the context of interaction.
-This perspective highlights the values about technical capabilities of AI models, including factuality (Roit et al., 2023; Hu et al., 2024), calibration (Liu et al., 2024a), output diversity (Bradley et al., 2024), and model inductive bias (Si et al., 2023).
-Additionally, this code includes research on aligning model behaviors with the characteristics and preferences of individual humans, e.g., predict human moral judgements and decisions (Jin et al., 2022), and cognitive biases (Jones and Steinhardt, 2022; Koo et al., 2023).
-Social sources mean values from the social groups include universal requirements for smooth functioning and survival of groups (Schwartz, 1994; Ammanabrolu et al., 2022).
-Value alignment at this level transcends personal interactions and emphasizes the broader categories defined by shared experiences, identities, cultures, norms, and more (Santy et al., 2023).
-Research in this area often targets the alignment of AI behaviors with the general preferences of humans (Hendrycks et al., 2021; Bakker et al., 2022; Lee et al., 2023a).
-Additionally, there are efforts aimed to align AI with specific targeted groups, examining issues through the lenses of fairness (Fleisig et al., 2023), social norms (Sky et al., 2023), morality (Prabhumoye et al., 2021; Rao et al., 2023), and beyond.
-Interactive sources considers values at the interaction level to include universal requisites of coordinated social interaction (Schwartz, 1994, 2012), which typically occurs in interpersonal situations, such as in the dynamics of speaker-receiver relationships during language communication (Hovy and Yang, 2021; Cabrera et al., 2023).
-In the context of “human-AI alignment”, we adjust the definition of Interaction values to be the AI values that humans expect for AI as tools, such as Usability (Zhang et al., 2023a), Autonomy (Wang et al., 2023a), among others.
-Research along this line focuses on alignment strategies to enhance human-AI interaction (Gupta et al., 2023; Oh et al., 2023), collaborative decision-making (Gajos and Mamykina, 2022; Yang et al., 2023a), and trust (Buçinca et al., 2021; Dhuliawala et al., 2023).
-Types of Values.
-In this dimension, we introduce the five high-order human values derived from a combination of the Schwartz Human Values (Schwartz, 1994) and empirical values studies from the surveyed papers.
-We provide a more in-depth taxonomy of value types in Section 5.2, including the relationship of these five types and more nuanced value categories.
-Self-Enhancement refers to a set of self-protective and personal values that emphasize enhancing self-esteem and a sense of personal worth (Sedikides and Strube, 1995; He et al., 2023; Xu et al., 2024; Huang et al., 2023; Park et al., 2023; Kang et al., 2023; Ye et al., 2024; Lima et al., 2021; Santurkar et al., 2023).
-One of the most important aspect is achievement, is competence as judged by social standards, which includes general capability (effectiveness/efficiency) as covered by the majority of the research (Guo et al., 2024).
-Another dimension is power, which relates to social status and prestige, as well as control or dominance over people and resources (Pinski et al., 2023).
-Self-Transcendence refers to a set of self-expanding and socially-focused human values that emphasize expanding beyond oneself (Frankl, 1966; Banovic et al., 2023; Wu et al., 2023d; Jin et al., 2022; Wu et al., 2023b; Huang et al., 2023; Ma et al., 2023b; Isajanyan et al., 2024; He et al., 2023; Nie et al., 2023; Santurkar et al., 2023; Ramezani and Xu, 2023; Sheng et al., 2021; Goldfarb-Tarrant et al., 2023).
-One critical aspect is benevolence, which relates to preservation and enhancement of the welfare of people with whom one is in frequent personal contact.
-Extensive efforts have investigated the values including helpfulness (Bai et al., 2022a), honesty/factuality (Hu et al., 2024), responsibility/accountability (Solyst et al., 2023).
-Another critical aspect is universalism, which relates to understanding, appreciation, tolerance and protection for the welfare of all people and for nature.
-Researchers have looked into values including inclusion/broad-mindedness (Bakker et al., 2022; Dong et al., 2024) and equality/fairness (Shen et al., 2022; Wang et al., 2023b).
-Conservation refers to a set of self-protective and socially-focused human values that hold and safeguard traditional institutions and customs (Hamilton, 2020; Qi et al., 2023; Zagalsky et al., 2021; Lai et al., 2022; Santy et al., 2023; Zhu et al., 2023; Zhou et al., 2023a).
-Under conservative values, people are concerned about security (Ssafety, harmony, and stability of society, relationships, and oneself), tradition (Respect for and acceptance of the customs and ideas provided by traditional culture or religion), and conformity (restraint of actions, inclinations, and impulses likely to upset or harm others and violate social expectations or norms).
-Exemplar studies in this field include safety (Anwar et al., 2024), mental health (Liu et al., 2023), and cultural moral norms (Ramezani and Xu, 2023).
-Openness to Change refers to a set of self-expanding and personally-focused human values motivated by an anxiety-free need to grow, in contrast to conservation (Bakker et al., 2022; Lima et al., 2020; Ashktorab et al., 2020; Jiang et al., 2023; Lima et al., 2020; Kang et al., 2023; Varanasi and Goyal, 2023; Yan et al., 2023; Hosking et al., 2024; Wu et al., 2023d; Kapania et al., 2023).
-Under this type, people are concerned about stimulation (seeking excitement, novelty, and challenges in life), hedonism (pursuing pleasure and sensuous gratification for oneself), and self-direction (independent thought and action—choices, creativity, and exploration).
-Exemplar studies in this field include understanding of privacy (Mireshghallah et al., 2024) and creativity (Isajanyan et al., 2024; Ashkinaze et al., 2024).
-We adopt the four high-order human values from Schwartz Human Values (Schwartz, 1994), supplementing them with additional values empirically collected from survey papers as shown in Figure 2.
-Additionally, as the conventional theory is missing the context of AI-system and human-AI interactions, we introduce a new high-order value type called Desired Values for AI Tools.
-This category encompasses the values humans expect from AI when used as tools in “human-AI interaction” (Yang et al., 2023b; Zhang et al., 2023a; Oh et al., 2023; Bakker et al., 2022; Boyd, 2021; Wang et al., 2023a; Paleja et al., 2021).
-These values include usability (competence according to the human experience on AI functionality) and human-likeness (resemblance to human intelligence and behavior).
-Exemplary studies in this area include assessments of AI usability (Zhang et al., 2023a) and autonomy (Wang et al., 2023a).
-4.1.2. Interaction Techniques to Specify AI Values. How humans could interactively specify values in AI development?
-This sub-research question investigates how human values are interactively111 We define ”interactions” broadly to encompass both ”synchronous” and ”asynchronous” interactions between humans and AI: (1) “Synchronous Interactions” indicates real-time exchanges where humans and AI systems interact simultaneously.
-(2) “Asynchronous Interactions” allows for delays between actions and responses, such as data annotations by humans to train the AI models.
-specified for AI systems to ensure alignment.
-It aims to elucidate the interaction techniques by which AI systems manifest or instantiate human values, thereby revealing the underlying mechanisms that shape their behavior or functionality.
-Explicit Human Feedback.
-This dimension refers to the direct specification of human values through explicitly defined formats or mechanisms.
-Principles provides AI systems with explicitly defined principles, guidelines, or rules that dictate behavior or decision-making in alignment with human values (Petridis et al., 2024a; Bai et al., 2022b).
-Rating and Ranking is widely used to assign numerical scores or rankings to options or outcomes based on their alignment with human values (Dai et al., 2024; Yuan et al., 2023).
-Natural Language Interaction/Conversations allows humans to interact with AI systems through natural language interfaces to express and communicate human values (Bai et al., 2022b; Hwang et al., 2023; Guo et al., 2024; Feng and Boyd-Graber, 2019; Yao et al., 2023a; Lin et al., 2024; Hong et al., 2022; Shen et al., 2023a; Zhang et al., 2021).
-For Multimodal Feedback, human values can also be provided in multiple modalities, such as sketches/images and gestures (Chung et al., 2022), to convey human values (Chen et al., 2022; Geva et al., 2022; Wu et al., 2022; Jörke et al., 2023).
-Implicit Human Feedback.
-This dimension refers to the indirect representation or inference of human values within AI systems through patterns, signals, or cues embedded in the data or decision-making processes.
-Discarded Options refers to the options or choices that human discard when interacting with AI systems during the decision-making processes, which also potentially infer human values (Park et al., 2021b).
-Language Analysis means that textual data and language patterns can also contain rich information to identify implicit references to human values or value-related concepts (Lu et al., 2023).
-Theory of Mind refers to the ability of agents and people to attribute mental states, such as beliefs, intentions, desires, emotions, knowledge, percepts, and non-literal communication, to themselves and others (Ma et al., 2023b).
-Social Relationships refers to the implicit values derived from analyzing human social relations and behaviors, which are derived from external sources (e.g., social network) that can inherently reflect their values (Isajanyan et al., 2024).
-Simulated Human Value Feedback.
-When human values in explicit formats are expensive or impossible to collect, one may simulate human-like feedback within AI systems to approximate human responses and preferences regarding specific values.
-Human Feedback Simulators uses computational algorithms to simulate human-like feedback on values, based on predefined criteria or training data (Petridis et al., 2024b; Li et al., 2023a).
-Comparison to Human Data refers to developing techniques that assess the likelihood or probability of AI-generated outputs matching human behaviors in a reference set or dataset (Dong et al., 2024).
-Synthetic Data curates data by generating synthetic comparisons based on naive assumptions or heuristic rules, followed by post-validation to ensure feedback quality (Kim et al., 2023).
-4.2. Align AI to Humans: [Uncaptioned image] Integrating Human Specifications into AI
-RQ2.
-How can human values be integrated into the development of AI?
-Existing studies have explored diverse methods to integrate human values into AI.
-We structure them by summarizing how to integrate general human values (Section 4.2.1) and customized human values (Section 4.2.2) throughout AI development stages?, and then elaborating what are the evaluation methods (Section 4.2.3) and supported platforms (Section 4.2.4) for the AI development?
-Additionally, we answer the four “Sub-Research Questions” by introducing the answer “Dimensions“ and providing “Codes“ associated with “Example Papers”.
-4.2.1. Integrating General Values to AI: how to incorporate general human values into AI development?
-This sub-research question focuses on the process of incorporating broad, universally recognized human values into the development of AI systems.
-The goal is to ensure that AI systems align with overarching ethical principles and societal norms, thereby promoting trust, acceptance, and responsible use.
-Instruction Data.
-This dimension refers to the types of data and processes used to provide guidance or direction to AI systems during their development.
-Human Annotation makes data with human-generated labels or annotations that indicate the presence or relevance of specific human values (Shen et al., 2023b; Pei et al., 2022).
-Human-AI CoAnnotation leverages both human expertise and AI capabilities to collaboratively annotate the data  (Sun et al., 2023; Li et al., 2023b).
-Simulated Human Data generates synthetic or simulated data that mimics human behaviors, preferences, or decision-making processes to provide training signals for AI systems (Dubois et al., 2023; Kim et al., 2023).
-Model Learning.
-This dimension refers to the model architecture design and training stages after the data collection, where human values are integrated during the model learning process.
-Online Human Alignment integrates human values into AI systems in real-time or during active system operation, often through interactive feedback loops or adaptive learning mechanisms.
-Examples include real-time user feedback and online training (Ouyang et al., 2022; Dong et al., 2023).
-Offline Human Alignment incorporates human values into AI systems prior to deployment or during offline training phases, without direct user interaction (Rafailov et al., 2023; Song et al., 2024; Yuan et al., 2023).
-Inference Stage.
-This dimension involves evaluating the alignment of AI systems with human values and assessing their performance and behavior in relation to predefined criteria or benchmarks.
-Prompting leverages prompting methods, such as in-context learning and chain-of-thought, on trained AI systems to elicit or critique AI regarding the encoded values (Bai et al., 2022b; Wang et al., 2023a).
-External Tool Interaction integrates external tools, like code interpreter for debugging, to cross-check and refine their initial generated content (Gou et al., 2024).
-Response Search generates a diverse range of high-quality outputs from which to choose (Bradley et al., 2024).
-4.2.2. Customizing AI Values: how to customize AI to incorporate values from individuals or human groups?
-This sub-research question explores the customization of AI values to align with specific contexts, domains, or user preferences.
-The goal is to enhance the alignment of AI systems within specific application domains or user communities.
-Customized Data.
-Finetuning with Curated Datasets aims to curate datasets for specific individuals or societal groups, and further finetune the pre-trained AI models on these specific datasetsto align them with targeted human groups and values (Felkner et al., 2023).
-These curated datasets include data collected from socio-demographic groups (Orlikowski et al., 2023), users’ history data (Lyu et al., 2023), expert-selected data for imitation learning (Cai et al., 2023) and others.
-Adapt Model by Learning.
-This dimension involves refining or adjusting AI values through techniques for customization, such as iterative learning process, model enhancements, or structural modifications.
-Group-based Learning trains AI models from specific user groups or communities to capture group-specific values or preferences (Zhao et al., 2023).
-Active Learning aims to interative selecting and lebeling data samples for AI model training based on their potential to improve alignment with user preference or values (Peschl et al., 2022).
-Inserted Adapters incorporates adapter modules or components into AI model architectures to fine-tune specific aspects or behaviors (Sun et al., 2022).
-Mixture of Experts combines multiple specialized models or experts to collectively capture diverse perspectives and values, with each expert focusing on a specific subset of the data or problem space (Shen et al., 2022).
-Enhanced Knowledge enhances AI model’s representations and embeddings with additional knowledge or context to improve alignment with user preferences or values (Wang et al., 2023c; Dalvi et al., 2022).
-Interactive Alignment This dimension involves actively engaging users or stakeholders in the process of customizing AI values to align with specific contexts, domains, or user preferences.
-Interactive Learning enables the users to provide feedback or corrections to AI models in real time, such as using interactive tutorials and user-driven customization interfaces (Maghakian et al., 2023; Swazinna et al., 2023).
-Steering Prompts provides users with prompts or cues to steer the behavior or decision-making of AI systems towards desired outcomes or values (Lahoti et al., 2023; Fei et al., 2023).
-Proactive Alignment anticipates user needs and preferences based on historical data or user profiles and proactively adjusting AI systems accordingly (Welch et al., 2022; Shen et al., 2023a).
-4.2.3. Evaluating AI Systems: how to evaluate AI regarding human values?
-The rise in the use of LLMs has also seen the rise of automatic evaluation of generated natural language text evaluation in different contexts.
-But in particular, a research dimension has focused on analyzing how closely values discussed in humane context has been adapted to AI models/applications and how these are being evaluated.
-Human-In-The-Loop Evaluation.
-This dimension involves incorporating human judgement, feedback, or interaction into the evaluation process to assess the effectiveness, robustness, and ethical implications of integrating human values into AI systems.
-Human Evaluation solicits feedback, opinions, or assessments from human evaluators to gauge the alignment of AI systems with human values and ethical standards (Liu et al., 2022; Sharma et al., 2023; Ouyang et al., 2023; Kiesel et al., 2022; Hosking et al., 2024).
-Human-AI Collaborative Evaluation collaboratively evaluates AI systems with both human evaluators and large language models (LLMs) to leverage the strengths of both human judgement and AI capabilities (Kwon and Mihindukulasooriya, 2023; Ammanabrolu et al., 2022).
-Automatic Evaluation.
-This dimension involves using computational methods or algorithms to assess the alignment of AI systems with human values, without direct human involvement.
-Human Simulators use simulation models or virtual agents to mimic human behavior and assess AI performance in human-like scenarios.
-Typical methods include agent-based simulations, synthetic user models, and others (Jia et al., 2024; Wang et al., 2023d; Felkner et al., 2023; Ruan et al., 2023).
-Evaluation Benchmarks establishes standardized benchmarks or metrics for evaluating AI performance in relation to human values and ethical considerations (Sun et al., 2023; Ramezani and Xu, 2023; Hu et al., 2024; Kiesel et al., 2022; Rao et al., 2023; Lee et al., 2023a; Sheng et al., 2021; Sharma et al., 2023).
-Distribution Difference compares the difference between the output distribution from AI generations and human data to evaluate AI (Patel and Pavlick, 2021; Lee et al., 2023a).
-Figure 3.
-The fine-grained topology of “Align Human to AI” direction, which studies humans’ cognitive and behavioral adaptation to the AI systems, which aims to help humans better critique, collaborate with, and co-adapt to AI.
-4.2.4. Ecosystem and Platforms: how to build the ecosystem to facilitate human-AI alignment?
+How can human values be integrated into the AI systems?
+  • RQ1: Human Values and Specifications.
+To identify key human values and specifications for AI alignment, we begin by addressing two critical subquestions: (1) what values have AI systems been aligned with?
+and (2) how can humans interactively specify values during AI development?
+As summarized at the top of Table 1, we present the key dimensions that emerged from our analysis of these questions, along with supporting studies.
+Categorizing Aligned Human Values.
+To systematically understand human values relevant to human-AI alignment shen2024valuecompass ; shen2025mind , we draw on the adapted Schwartz Theory of Basic Values, examining values along two dimensions: Sources and Types.
+The Sources dimension categorizes values as individual (e.g., personal interests and biological needs like factuality or cognitive biases) schwartz1994there ; kiesel2022identifying ; jin2022make , social (e.g., shared group norms such as fairness or morality) schwartz1994there ; ammanabrolu2022aligning ; santy2023nlpositionality , and interactive (e.g., expectations in human-AI interactions like usability, autonomy, and trust) hovy2021importance ; cabrera2023zeno ; wang2023humanoid .
+The Types dimension organizes values into four high-order categories: Self-Enhancement (e.g., achievement, power) sedikides1995multiply ; he2023interaction ; xu2023lemur , Self-Transcendence (e.g., benevolence, honesty, fairness) frankl1966self ; banovic2023being ; wu2023cross , Conservation (e.g., safety, tradition, conformity) conservatism ; qi2023fine ; zagalsky2021design , Openness to Change (e.g., creativity, privacy, autonomy) bakker2022fine ; lima2020collecting ; ashktorab2020human .
+These dimensions offer a comprehensive framework for evaluating and aligning AI systems with the multifaceted nature of human values.
+Interaction Techniques to Specify AI Values.
+This sub-research question explores how human values are interactively specified to ensure AI alignment, focusing on the techniques through which AI systems manifest or internalize these values.
+It identifies three main approaches: explicit human feedback, where values are directly communicated via principles, ratings, natural language interactions, or multimodal inputs like gestures and images petridis2024constitutionmaker ; dai2023safe ; bai2022constitutional ; implicit human feedback, where values are inferred from indirect cues such as discarded options, language patterns, theory of mind reasoning, and social relationships isajanyan2024social ; park2020scalable ; lu2022towards ; and simulated human value feedback, where AI systems approximate human responses using feedback simulators, comparisons to human data, or synthetically generated data petridis2024constitutionalexperts ; dong2023aligndiff ; kim2023aligning .
+Together, these approaches illuminate the mechanisms by which AI systems interpret and enact human values via direct and indirect human-AI interaction.
+Key Takeaways.
+By comparing prior research with our comprehensive analysis of human values and interaction techniques, we found that existing studies are largely constrained to conventional principles and standard interaction methods, overlooking the broader spectrum of human values and the interactive approaches needed to specify them effectively in alignment.
+  • RQ2: Integrating Human Specifications into AI.
+Building on the value-laden human specifications gathered through interaction, we next explore diverse methods for integrating human values into AI systems.
+We examine this central question across two key stages of the AI lifecycle—development and deployment—by asking: (1) how can general or customized human values be integrated throughout the AI development process?
+and (2) what methods and platforms are available to evaluate values during AI development?
+Integrating General Values to AI.
+This sub-research question examines how broad, universally recognized human values are embedded into AI systems to ensure ethical alignment and societal acceptance.
+It outlines three key dimensions: Instruction Data, which includes human annotations, human-AI co-annotation, and simulated human data to guide value-based training shen2023multiturncleanup ; sun2024principle ; dubois2024alpacafarm ; Model Learning, where human values are integrated during training through either real-time online alignment or offline processes prior to deployment ouyang2022training ; dong2023raft ; rafailov2024direct ; and Inference Stage, where AI systems are evaluated and refined using techniques such as prompting, external tool interactions, and response search to ensure their outputs align with predefined ethical criteria bai2022constitutional ; wang2023humanoid ; gou2023critic ; bradley2023quality .
+Together, these processes aim to build AI systems that promote trust and responsible use.
+Customizing AI Values.
+This sub-research question investigates how AI systems can be customized to reflect specific user preferences, application domains, or community values, thereby improving contextual alignment.
+It identifies three primary strategies: Customized Data, which involves curating and finetuning datasets based on socio-demographic groups, user histories, or expert selections to align models with targeted human values felkner2023winoqueer ; orlikowski2023ecological ; lyu2023exploiting ; Adapt Model by Learning, which includes techniques such as group-based learning, active learning, adapter insertion, mixture of experts, and enhanced knowledge integration to refine model behavior zhao2023group ; peschl2022moral ; sun2022bertscore ; and Interactive Alignment, which actively engages users through real-time feedback, steering prompts, and proactive adjustments based on user profiles to tailor AI systems to specific contexts and preferences maghakian2022personalized ; swazinna2022user ; lahoti2023improving .
+Evaluating AI Systems.
+This sub-research question examines how the integration of human values into AI systems, particularly large language models (LLMs), is evaluated, highlighting both human-in-the-loop and automated methods.
+Human-in-the-loop evaluation involves human judgment, feedback, and collaboration to assess the ethical and value alignment of AI outputs, leveraging direct human input or combined human-AI assessment processes sharma2023cognitive ; ouyang2023shifted ; ammanabrolu2022aligning .
+Automatic evaluation utilizes computational techniques, including human simulators, standardized benchmarks, and distributional comparisons, to evaluate alignment without human intervention jia2024towards ; sun2024principle ; patel2021stated .
+Together, these approaches aim to ensure that AI systems reflect human ethical standards and value frameworks effectively and reliably.
+Ecosystem and Platforms.
 The ecosystem and platforms refer to the broader context in which AI systems operate and interact with other agents, platforms, or environments.
 This includes the infrastructure, frameworks, and technologies that support the development, deployment, and utilization of AI systems.
-LLM-based Agents are based on large language models (LLMs) such as GPT (Generative Pre-trained Transformer) models, which have been pre-trained on vast amounts of text data (Dubois et al., 2023; Xu et al., 2023; Wu et al., 2023a; Zhou et al., 2023b).
-RL-based Agents are based on reinforcement learning (RL) algorithms to learn and adapt their behavior based on feedback from the environment or human users (Yuan et al., 2024).
-Annotation Platforms refers to the ecosystems that are designed to crowdsource human demonstrations as collected data for reinforcement learning (Gerstgrasser et al., 2021) and supervised finetuning learning for alignment (Pei et al., 2022).
-    Underexplored Dimensions in Aligning AI to Humans (See Sec 5.1 for supporting data and evidence): 1.
-Implicit human feedback and simulated human value feedback are under-explored in existing research work.
-2.
-Developing and customizing AI during the inference stage or in an interactive way is under-explored.
-3.
-Human-in-the-loop evaluation is much less explored than automatic evaluation.
-DIRECTION-II: [Uncaptioned image] ALIGN HUMANS to AI
-[…]
-When interacting with people, AI agents do not just influence the state of the world – they also influence the actions people take in response to the agent, and even their underlying intentions and strategies (Hong et al., 2023).
-From a long-term perspective, it is essential to consider the dynamic changes around human-AI alignment.
-This direction outlines alignment research from the Human-centered perspective (e.g., HCI/Social Science domains) and provides HCI researchers and user experience designers with guidance for approaching two core research questions: RQ3.
-Human Cognitive Adjustment to AI (Section 4.3) and RQ4.
-Human Adaptive Behavior to AI (Section 4.4).
-4.3. Align Humans to AI: [Uncaptioned image] Human’s Perceptual Adaptation to AI
-RQ3.
-How might humans learn to perceive, explain, and critique AI?
-Humans needs to understand AI to better specify their demands and collaborate with AI.
-Also, as AI systems produce a range of risks, it is important to elicit humans’ critical thinking of AI instead of relying on AI blindly.
-Therefore, we categorize existing literature to answer the questions of how humans learn to perceive and understand AI (Section 4.3.1), and how to engage in critical thinking on AI?
-(Section 4.3.2).
-We also answer the two “Sub-Research Questions” by introducing the answer “Dimensions“ and providing “Codes“ associated with “Example Papers”.
-4.3.1. Perceiving and Understanding AI: how do humans learn to perceive and explain AI systems?
-Addressing the problem of human perception and understanding of AI includes the fundamental education and training required for less technical people to improve understanding of the behind mechanisms and outputs produced by AI systems.
-It includes visualizations and human-centered explanations tohelp more people learn to understand AI outputs.
-Educating and Training Humans.
-AI is increasingly incorporated into the daily lives of a broad spectrum of users, including those with little to no technical knowledge on how AI operates.
-Here we discuss how to help these less technical people become more AI literate, or better trained to use AI.
-AI Literacy and Awareness is broadly defined as the core competencies required for less technical people to better use and collaborate with AI (Long and Magerko, 2020).
-Beyond, there is also more in depth and explicit Training Courses to support individuals to better collaborate and utilize AI (Paranjape et al., 2019; McDonald and Pan, 2020).
-AI Sensemaking and Explanations.
-To improve people’s understanding of AI systems and generated outputs, various techniques and approaches have been developed to help people learn to make sense of and explain the model’s outputs.
-Perception and Sensemaking on AI involves the process through which humans learn to make better sense of AI mechanisms and decision makings (Kaur et al., 2022; Sivaraman et al., 2023; Petridis et al., 2023).
-People also need to understand how AI systems arrive at specific generated outputs.
-Prior studies in Human-Centered Explanations have examined various approaches and interactive techniques to increase human understanding of AI generations and outputs (Tenney et al., 2020; Tintarev and Masthoff, 2007; Smith-Renner et al., 2020).
-4.3.2. Critical Thinking around AI: how do humans think critically about AI systems?
-Beyond simply perceiving and understanding AI, individuals need to compare their mental model of AI with their own mental model to judge whether the AI is behaving rationally and ethically.
-Broadly, this involves exploring the ways in which humans engage in critical thinking and reflect on their interactions with and evolving understanding of AI technologies.
-Here we discuss studies that help humans become more capable to identify biases and errors in AI output, and the ethical implications that arise from using AI algorithms in decision-making processes.
-Also, we discuss how humans need to calibrate their mental models to be more aligned with AI.
-Trust and Reliance on AI Decisions refers to the extent to which people mentally trust the AI competency and practically rely on AI for output generation and decision making.
-We use the term Trustworthiness to indicate whether humans decide to trust the reliability, integrity, and competence of an AI system in delivering accurate and trustworthy decisions or recommendations to the users (Yin et al., 2019; Ma et al., 2023a; Srivastava et al., 2023).
-We define AI Reliability as to what extent humans utilize and rely on AI to automate decision making in practice with low error rates and robustness performance (Schemmer et al., 2023).
-(Schemmer et al., 2023; Vasconcelos et al., 2023; Suresh et al., 2022).
-Selective AI Adoption refers to the criteria and considerations that guide the adoption or rejection of AI technologies based on their perceived benefits, risks, and alignment with user needs and preferences (Gu et al., 2024; Vasconcelos et al., 2023).
-Ethical Considerations and AI Auditing refers to potential moral and societal issues that arise from the development, deployment and use of AI systems, as well as systematic examination and evaluation of AI systems regarding these issues.
-Ethical Concerns and AI Mitigation encourages humans to carefully consider if the AI systems possess ethical concerns (e.g., such as bias and discrimination, privacy violations and the potential for harm or misuse) and implement strategies and practices to address and reduce these ethical concerns associated with AI (Holstein et al., 2019).
-Audit Deployed AI refers to the systematic examination of AI systems to ensure that they operate as intended, comply with ethical and legal standards, and do not cause unintended harm (Bandy, 2021).
-Re-Calibrating Cognition to Align with AI.
-This sub-category deals with interventions and techniques to help users adjust (1) their own mental model of how the AI operates and (2) recalibrate their perception of the AI.
-Updating Human Mental Model of AI means that when the AI is not aligned with humans, it is important to adjust human perceptions of the confidence in AI systems based on the model’s capability and track record (He et al., 2023; Kocielnik et al., 2019) .
-Recalibrating Trust and Reliability indicates that humans adjust their perceptions of trust and reliability in AI systems based on their performance and reliability.
-This is also important to foster appropriate reliance and skepticism (Wischnewski et al., 2023).
-4.4. Align Humans to AI: [Uncaptioned image] Human’s Behavioral Adaptation to AI
-RQ4.
-How do humans and society make behavioral changes and react to AI advancement?
-As AI becomes increasingly integrated into daily life, it is essential to understand its influence on humans, encompassing both positive and negative aspects.
-Moreover, it is crucial to determine how individuals and society can best and most appropriately respond to this influence.
-To this end, we summarize literature to answer how do humans learn to collaborate with AI in diverse AI roles?
-(Section 4.4.1), how humans and society are impacted by AI (Section 4.4.2) and how might we assess these impacts?
-(Section 4.4.3) We articulate the three “Sub-Research Questions” by introducing the answer “Dimensions“ and providing “Codes“ associated with “Example Papers”.
-4.4.1. Human-AI Collaboration Mechanisms: what are human strategies to collaborate with AI that have differing levels of capabilities?
-This category looks at many ways that humans and AI can collaborate, such as teamwork, co-creation, and coproduction.
-AI Assistant for Humans captures the essence of a symbiotic relationship where AI systems are designed to bolster human capabilities, with humans steering the interactions.
-LLM-based Demand Specification employs LLMs to interpret and respond to human requests, powering virtual assistants and chatbots that streamline information retrieval and improve task accuracy (Terry et al., 2023).
-This naturally extends to LLM-based Prompt Strategy, which helps humans to better write prompts using additional abstraction and scaffolding methods.
-The example system can enhance humans’ capability in generating intelligent prompts and suggestions, such as autocomplete and question-generation tools for LLM-based systems (Wu et al., 2022; Wang et al., 2024), facilitating humans’ smooth and intuitive decision-making processes.
-In the creative arena, LLM-based Prototypes utilize AI to transform human ideas into tangible or organized concepts  (Gebreegziabher et al., 2023), enabling professionals to explore and refine a wide array of artistic possibilities with AI-generated options (Huh et al., 2023).
-On a broader note, researchers have also explored how AI-Assisted Task and Decisions aims to achieve complementary performance for human-AI collaborative tasks by empower humans to discern when and how to adopt AI-assisted recommendations or AI-generated explanations for decision-making (Choi et al., 2019; Wang et al., 2022; Wang and Yin, 2023; Zheng et al., 2023; Wu et al., 2023c).
-Human-AI Partnership refers to a collaborative relationship where humans and AI systems work together as partners, combining their respective strengths to achieve shared goals more effectively.
-Simulated Agency enables humans to collaborate with AI partners with simulated agency or autonomy (e.g., autonomous agents and collaborative robots (Park et al., 2023)) to make decisions collaboratively.
-Reciprocal Learning focuses on how humans learn from and exchange knowledge with AI systems, enhancing human-AI collective capabilities and performance through knowledge-sharing platforms and collaborative filtering (Zagalsky et al., 2021; Pinski et al., 2023).
-AI Delegation allows humans to delegate AI partners to help finish tasks or responsibilities, facilitated by task assignment algorithms and workflow management systems (Pinski et al., 2023).
-Furthermore, AI-Mediated Tasks emphasizes how traditional human tasks or behaviors (e.g., communication) would be changed by incorporating AI partners in the loop (Bernstein et al., 2010; Sundar and Lee, 2022; Xu and Zhang, 2023).
-Co-Design with AI explores AI as a design material or a partner in prototyping, which enables humans to converse with AI in situations that can collaboratively improve the design outcomes.
-(Yildirim et al., 2022; Wang et al., 2024)
-AI Tutoring for Human Learning investigates how humans improve their learning and knowledge throughinteractions with AI tutors that can perform better than humans in some tasks.
-With the tailored instructions and customized feedback from AI tutors, humans can enhance their the learning outcomes and facilitate mastery of new skills more effectively than traditional learning methods.
-AI Tutor for Technical Skills refers to empowering humans to learn technical skills from AI tutors.
-For technical subjects like coding, AI tutors can analyze a learner’s progress in real time, adjusting the pace, content, and approach to ensure a solid understanding of complex concepts and practical skills such as programming (Ma et al., 2024).
-Similarly, AI Tutor for Social and Behavioral Skills involves enabling humans to learn social and behavioral skills from AI tutors.
-Humans can leverage virtual simulations, created by AI tutoring systems, to practice public speaking, interpersonal communication, and other soft skills.
-By analyzing verbal and non-verbal cues, humans can receive constructive feedback on areas such as body language, tone, and delivery from AI, ultimately enhancing their ability to communicate effectively across various settings (Peng et al., 2021; Shaikh et al., 2024).
-4.4.2. AI Impact on Humans and Society: how are humans influenced by AI systems ?
-This category explores the effects of AI advancement on human behaviors, attitudes, and societal dynamics.
-It involves examining the behavioral changes, adaptations, and reactions that individuals, groups and wider communities undergo in response to the proliferation of AI technologies.
-The goal is to elucidate the multifaceted impacts of AI on human behavior and society and to inform policy-making, education, and intervention efforts.
-Impacts on Participatory Individuals and Groups covers the effects of AI advancement on the behaviors, attitudes, and experiences of both individuals and groups.
-This dimension focuses on examining how AI technologies influence decision-making, creativity, privacy, and authorship.
-Decision Making refers to analyzing how AI technologies influence human decision-making processes, including biases, preferences, and risk assessment, in various domains such as healthcare, finance, and personal life (Shen and Huang, 2020).
-Human Creativity explores the impact of AI technologies on human creativity, innovation, and expression, including the augmentation or automation of creative tasks and the emergence of new forms of artistic expression (Ashkinaze et al., 2024).
-Privacy relates to investigating the implications of AI technologies for individual privacy rights, data protection, and surveillance, including concerns about data collection, tracking, and algorithmic profiling (Li et al., 2024a).
-Authorship catalogs issues related to intellectual property, attribution, and ownership of AI-generated content, including questions of legal responsibility, copyright infringement, and plagiarism detection.
-Salient is that AI can produce increasingly realistic, synthetic data quickly and at low cost, which brings forth tensions around the use of such data to make decisions (Hämäläinen et al., 2023; Atreja et al., 2023).
-Societal Concerns and AI Impacts involves the broader societal implications and consequences of AI advancement on misinformation, education, social relationships, norms, job displacement, and other aspects of human society.
-Misinformation and Moderation concerns the challenges of misinformation, disinformation, and online content moderation in the context of AI-driven information ecosystems, including concerns about algorithmic bias and filter bubbles (Atreja et al., 2023).
-Impacts on Education pertains to assessing the effects of AI technologies on education systems, learning outcomes, pedagogical practices, and workforce training, including opportunities for personalized learning and skill development (Kazemitabaar et al., 2023).
-Impacts on Social Relationship and Norms explores how AI technologies shape interpersonal relationships, social interactions, and cultural norms, including changes in communication patterns, social dynamics, and ethical considerations (Cheng et al., 2022).
-Workplace refers to examining the effects of automation and AI technologies on employment patterns, job markets, and workforce dynamics, including concerns about job displacement, re-skilling, and economic inequality (Park et al., 2021a).
-Reaction to AI Advancement involves societal responses, regulatory frameworks, and policy initiatives aimed at addressing the challenges and opportunities posed by AI technologies.
-This dimension encompasses efforts to regulate AI deployment, re-calibrate societal acceptance, and manage potential backlash.
-For example, reaction to bias and discrimination in algorithmic decision making can depend on how people perceive the machine and the context of use, i.e., if the machine is considered an actor embedded in social structures that call for blame when harmful decisions are made (Lima et al., 2023a).
-AI Regulatory and Policy includes regulatory frameworks, legal frameworks, and policy initiatives aimed at governing AI development, deployment, and use, including concerns about ethics, safety, and accountability (Hacker et al., 2023; Lucaj et al., 2023).
-Shifts in AI Acceptance relates to investigating societal attitudes, organizational practices, and acceptance of AI technologies over time in practice, including shifts in public opinion and AI utilization by humans and institutes regarding AI deployment and impact (Lima et al., 2023b; Pruss, 2023).
-Surveillance and Intervention emphasizes the need for transparency, monitoring, and human oversight in the algorithmic decision-making process.
-This approach enables better human control over AI systems and helps mitigate potential risks associated with their use (Park et al., 2021a).
-Responsible AI Checklists involves the creation of ethical guidelines, such as those focusing on fairness and transparency, to ensure the responsible development and deployment of AI systems.
-These published principles serve as a foundation for guiding ethical AI practices (Madaio et al., 2020).
-4.4.3. Evaluation in Human Studies: how might we evaluate and understand the impact of AI on humans and society?
-We summarize common empirical methods used to rigorously understand and assess the impact of AI on humans.
-Specifically, we focus on two types of impact.
-On the micro-level, we discuss how to evaluate the effectiveness of human-AI collaboration; on the macro-level, we discuss how to assess the impact of AI on a large group of people over a long period of time.
-Evaluate Human-AI Collaboration refers to the evaluation of the effectiveness of an AI system in collaboration with humans.
-It is key to not only consider the final output, but also the interaction experience (Lee et al., 2023b).
-Human-AI Team Performance compares human-AI team performance with the performance of humans alone without AI collaboration The metrics for measuring performance should include both task success metrics (e.g., accuracy) as well as indicators of efficiency (Mozannar et al., 2024; Ding et al., 2023a).
-Cognitive Workload and User Satisfaction involves understanding the degree of cognitive load that the user experiences when interacting with the system, as well as user satisfaction with the interaction and the final outcome.
-Such aspects are often captured via surveys or interviews (Ding et al., 2023a; Lee et al., 2023b).
-Control and Trust refers to how user control can support the avoidance of catastrophic AI failures, especially in high-stakes settings where AI mistakes could lead to harm (Srikanth et al., 2024; Ding et al., 2023b).
-Interaction Behavioral Analytics refers to measuring task performance quantitatively.
-This approach includes recording user interaction data and analyzing the patterns (Villalba et al., 2023; Perry et al., 2023).
-Qualitative Survey and Interviews refers to qualitative approaches to understanding human-AI interaction.
-Commonly used methods include qualitative survey questions (i.e., open-ended) and user interviews to assess aspects of the user experience.
-(Zhang et al., 2023b).
-Statistical Analysis utilizes methods such as regression analysis to quantitatively analyze and evaluate data from human studies, allowing for the verification of hypotheses (Wang et al., 2021).
-Evaluate Societal Impact refers to the macro-impact of a group of people as they come to use AI broadly.
-This dimension requires sufficient scale and time.
-The aim is to understand how the group’s behavior changes as people within it frequently interact with AI.
-Public Opinion Surveys aims to investigate the impact of AI on human measures of interest though deploying and analyzing large scale questionnaires  (Santurkar et al., 2023).
-Behavioral Data Analytics collects large-scale and potentially longitudinal behavior data, with the aim of understanding how pattern evolve over time (Shin et al., 2023; Zhou and Lee, 2024).
-    Underexplored Dimensions in Aligning Humans to AI (See Sec 5.1 for supporting data and evidence): 1.
-Educating and training humans on AI literacy is under-investigated.
-2.
-Auditing AI for various ethical values is not fully explored.
-3.
-The collaboration between humans and AI with similar or superior capabilities is under-explored.
-4.
-The societal impacts of and reactions to AI advancements are not fully explored.
-5. Findings and Discussions on Current Gaps
-In this section, our aim is to consolidate key findings derived from our analysis of the framework and the current state of literature we reviewed.
-We begin by analyzing the overall trends and gaps in the literature (Section 5.1).
-We then focus on three essential aspects: the relationship between human values and alignment (Section 5.2), the potential interaction modes used to specify human values (Section 5.3).
-Figure 4.
-The number of papers for each dimension in the bidirectional human-AI alignment framework.
+LLM-based Agents are based on large language models (LLMs) such as GPT (Generative Pre-trained Transformer) models, which have been pre-trained on vast amounts of text data dubois2024alpacafarm ; xu2023gentopia ; wu2023autogen ; zhou2023sotopia .
+RL-based Agents are based on reinforcement learning (RL) algorithms to learn and adapt their behavior based on feedback from the environment or human users yuan2024uni .
+Annotation Platforms refers to the ecosystems that are designed to crowdsource human demonstrations as collected data for reinforcement learning gerstgrasser2021crowdplay and supervised finetuning learning for alignment pei2022potato .
+Key Takeaways.
+Our systematic analysis of value integration and evaluation in AI reveals a strong focus on explicit value annotations, while implicit value expressions and behaviors are often neglected.
+Despite the power of general-purpose AI, methods for customizing systems to reflect individual or group values remain underexplored.
+Additionally, there is a lack of standardized criteria for evaluating human-in-the-loop methods and supporting platforms, underscoring the need for more robust and context-sensitive evaluation frameworks in value-aligned AI development.
+3.2 Align Humans to AI
+From a long-term perspective, it is crucial to consider the dynamic and evolving nature of human-AI alignment.
+This direction emphasizes a human-centered perspective—drawing from fields such as HCI and the social sciences—and offers guidance for researchers and user experience designers in addressing two core research questions: RQ3.
+How can humans learn to perceive, explain, and critique AI?
+and RQ4.
+How do individuals and society adapt their behaviors in response to AI advancements?
+  • RQ3: Human Cognitive Adjustment to AI.
+For effective collaboration and value specification, humans need to develop a clear understanding of how AI systems function.
+As AI introduces various risks, fostering critical thinking is essential to prevent blind reliance.
+To address this, we systematically investigate, as summarized in Table 1, (1) how humans learn to perceive and understand AI, and (2) how they can engage in critical reflection on AI behavior and outputs.
+Perceiving and Understanding AI.
+This sub-research question explores how to enhance human understanding and perception of AI systems, particularly among non-technical users, through education, training, and human-centered explanation techniques.
+It emphasizes the importance of AI literacy and awareness as foundational competencies for effective human-AI collaboration long2020ai , supported by explicit training courses designed to improve users’ ability to engage with AI ketan-ai-medicine-education ; mcdonald2020intersectional .
+Additionally, it highlights efforts in AI sensemaking and human-centered explanations, including visualizations and interactive techniques, to help people interpret AI mechanisms and outputs, thereby fostering more informed and meaningful human-AI interactions kaur2022sensible ; sivaraman2023ignore ; tenney2020language .
+Critical Thinking around AI.
+This sub-research question examines how individuals critically reflect on and evaluate AI systems by comparing their own mental models with those of the AI, focusing on the rationality, reliability, and ethical behavior of these technologies.
+It emphasizes the need for humans to develop critical thinking skills to identify biases, errors, and ethical concerns in AI outputs, and to audit AI systems for compliance with moral and societal standards.
+Key areas include building appropriate levels of Trust and Reliance on AI based on its competence and reliability yin2019understanding ; ma2023should ; schemmer2023appropriate ; clark2025epistemic , engaging in selective AI adoption aligned with user needs and values gu2024ai ; vasconcelos2023explanations , and addressing Ethical Considerations through mitigation strategies and auditing holstein2019improving ; bandy2021problematic .
+Additionally, it underscores the importance of Recalibrating Cognition, where users adjust their understanding and expectations of AI performance and reliability to foster a balanced and informed relationship with AI systems he2023interaction ; kocielnik2019will ; wischnewski2023measuring .
+Key Takeaways.
+Our systematic review reveals a significant gap in research on training and educating humans to develop appropriate knowledge, trust, and reliance when collaborating with AI systems.
+Additionally, there is limited exploration of how to critically evaluate AI across a broad spectrum of human values.
+  • RQ4: Human Adaptation Behavior to AI.
+Building on our investigation of human cognitive adjustments to AI, we further examine how individuals and society can respond effectively to AI’s expanding influence.
+To guide this inquiry, we address three key questions: (1) How do humans learn to collaborate with AI across its diverse roles?
+(2) In what ways are individuals and society affected by AI?
+and (3) How can these impacts be comprehensively assessed?
+The key dimensions emerging from this analysis are summarized at the bottom of Table 1.
+Human-AI Collaboration Mechanisms.
+This category explores the diverse ways humans and AI collaborate through partnerships, co-creation, and mutual learning.
+AI Assistants, particularly those powered by LLMs, support human tasks by interpreting user demands, enhancing prompt formulation, generating creative prototypes, and aiding decision-making wu2022ai ; gebreegziabher2023patat ; scattershot .
+In collaborative frameworks, humans and AI function as AI Partners, where simulated agency and reciprocal learning enable joint decision-making and knowledge sharing park2023generative ; zagalsky2021design ; bernstein2010soylent .
+AI delegation and mediation transform traditional human tasks, while co-design with AI treats the system as both a collaborator and a design material.
+Additionally, AI Tutoring systems enhance human learning in both technical and social domains by offering tailored feedback, adaptive instruction, and immersive practice environments, ultimately improving skill acquisition and performance ma2023hypocompass ; peng2021say ; shaikh2023rehearsal .
+AI Impact on Humans and Society.
+This category investigates the multifaceted effects of AI advancement on human behavior, attitudes, and societal dynamics, aiming to inform policy, education, and interventions.
+The Impacts on Participatory Individuals and Groups dimension focuses on how AI influences human decision-making, creativity, privacy, and authorship—shaping behaviors and raising concerns about data rights and intellectual ownership shen2020useful ; ashkinaze2024ai ; li2024human ; shi2025siren .
+The Societal Concerns and AI Impacts dimension expands this lens to the societal level, examining AI’s effects on misinformation, education, social norms, and the workplace, including issues like disinformation, shifts in learning practices, changes in interpersonal relationships, and job displacement 10.1145/3610082 ; kazemitabaar2023novices ; cheng2022human .
+The Reaction to AI Advancement dimension explores regulatory, cultural, and institutional responses to AI, addressing how societies perceive, govern, and adapt to AI technologies lima2023blaming ; hacker2023regulating ; 10.1145/3593013.3593992 .
+This includes efforts to regulate bias and discrimination, develop policy frameworks, track evolving AI acceptance, ensure transparency and oversight, and establish responsible AI checklists to guide ethical and safe deployment.
+Evaluation in Human Studies.
+This summary covers common empirical methods used to rigorously evaluate AI’s impact on humans at both micro and macro levels.
+At the micro-level, Human-AI Collaboration Evaluation assesses not only task success and efficiency but also user experience, including cognitive workload, user satisfaction, control, and trust, especially in critical settings to prevent failures Lee2022EvaluatingHM ; Mozannar2024TheRE ; ding-etal-2023-harnessing ; Lee2022EvaluatingHM .
+Methods include quantitative interaction analytics, qualitative surveys and interviews, and statistical analyses to understand and verify user behaviors and perceptions.
+At the macro-level, Societal Impact Evaluation focuses on understanding long-term behavioral changes within large populations as AI use becomes widespread, employing large-scale public opinion surveys and behavioral data analytics over time to capture evolving patterns and societal shifts related to AI interaction santurkar2023whose ; Shin2023SuperhumanAI ; Zhou2024GenerativeAI .
+Key Takeaways.
+Our review reveals that while prior studies have extensively examined how AI can assist humans across various tasks, there is limited exploration of the challenges humans face when collaborating with AI that surpass human capabilities in certain domains.
+Additionally, more research is needed to understand the evolving impact of AI on individuals and society at large over time.
+4 Underexplored Research Gaps and Challenges
+Figure 2: The number of papers for each dimension in the bidirectional human-AI alignment framework.
 Out of papers that are relevant to each research question (i.e., gray bars), we show the number of papers that are relevant to each dimension (i.e., color bars).
-This figure illustrates the extent to which each dimension has been explored by existing research.
-We provide more analysis in Section 5.1.
-5.1. Meta Analysis of Trends and Gaps
-Based on our coding of all papers, we computed the number of relevant papers for each dimension in the proposed bi-directional framework.
-Based on the distribution in Figure 4, we noticed that in existing literature certain dimensions are over- or under-represented.
-We outline the under-explored dimensions in two directions in the highlighted grey bars in the end of Section 4.2 and Section 4.4, respectively, and provide more details below.
-Underexplored dimensions in AI-centered alignment research.
-Most literature specified human values using explicit human feedback, whereas implicit and simulated human feedback were largely under-explored.
-In the integrating human specifications to AI research question, Besides, developing and customizing AI models during the inference stage or using the interactive approach is under-explored.
-Also, human-in-the-loop evaluation is much less explored than automatic evaluation.
-Underexplored dimensions in Human-centered alignment research.
-Additionally, many studies on human cognitive adjustment to AI focus on enabling AI sensemaking and explanations so that humans can better understand, trust, and rely on AI.
-However, these studies often focused on explaining AI decision-making justification rather than educating people to acquire general skills and competencies required to understand, use, critique, and interact effectively with AI systems (i.e., AI literacy).
-AI literacy (Long and Magerko, 2020) plays a fundamental role in ensuring people understand and use AI correctly, and warrants deeper exploration.
-Also, we observed a wide range of studies (Wu et al., 2022; Gebreegziabher et al., 2023; Park et al., 2023; Ma et al., 2024) which developed interactive mechanisms and prototypes to empower human collaboration with AI.
-However, most studies assumed that AI plays an assistant role, being less capable than humans.
-This situation might change over time.
-Moreover, the influence of AI advancements on human behavior, social relationships, and societal changes is essential but remains largely unexplored.
-Figure 5.
-The interaction techniques for specifying values in human-AI alignment.
-We compare the common interaction techniques used for the model “Learning” and “Inference” stages in human-focused (e.g., HCI) and AI-focused (e.g., NLP/ML) research studies.
-    Takeaways of Interaction Techniques for Alignment.
-1.
-Some common human feedback formats (rating, ranking) used in NLP/ML are not often studied in HCI.
-2.
-Diverse human interactive feedbacks in HCI are not fully used in AI development in NLP/ML fields.
-5.2. Insights into Human Values for Alignment
-Our analysis, based on the adaptation of Schwartz’s Theory of Basic Values and our comprehensive literature review, identifies three critical findings for future research:
-[…]
-Universal vs.
-Personalized AI Values.
-While certain values are universally expected from AI (e.g., capability, equity, responsibility), others may be undesirable in specific contexts (Ngo et al., 2024) (e.g., seeking power).
-Simultaneously, AI models should be adaptable to diverse human value systems (Sorensen et al., 2024).
-Research is needed to develop methods for identifying appropriate value sets for specific individuals or groups, and for customizing AI to align with user values while maintaining ethical principles.
-Disparities in Value Expectations and Evaluation.
-The fundamental differences between humans and AI necessitate distinct approaches to value evaluation.
-For instance, assessing AI honesty may require mechanistic interpretability (Bereska and Gavves, 2024), a more rigorous standard than that applied to humans.
-Future studies should explore methods for evaluating and explaining AI values and calibrating human expectations accordingly.
-5.3. Interaction Techniques for Specifying Human Values: A Cross-Domain Analysis
-Our research reveals disparities in interaction techniques for human-AI value alignment across AI-centered (NLP/ML) and human-centered (HCI) domains.
-As depicted in Figure 5, this analysis focuses on three key areas:
-Domain-Specific Interaction Techniques.
-The interaction techniques in AI-centered (NLP/ML) and Human-centered (HCI) alignment studies are often differ (Brath, 2021).
-NLP/ML studies primarily utilize numeric and natural language-based techniques.
-Also, NLP/ML research explore implicit feedback to extract human hidden feedback.
-In contrast, HCI research encompasses a broader range of graphical and multi-modal interaction signals (e.g., sketches, location information) beyond text and images.
-This disparity suggests potential gaps in extracting comprehensive human behavioral information.
-Stage-Specific Interaction Techniques.
-In NLP/ML, the learning stage predominantly employs rating and ranking interactions for alignment in dataset generation.
-However, when humans use AI in the inference stage, as demonstrated in HCI research, involves more diverse user interactions.
-This discrepancy highlights the need for alignment between model development and practical deployment.
-Divergent Data Utilization.
-NLP/ML typically uses interaction outputs as training datasets, while HCI analyzes this data to understand human behavior and feedback.
-As AI systems evolve, developing new interaction modes to capture a broader spectrum of human expression becomes crucial.
-6. Future Directions
-Drawing upon insights gained from the development of our framework and the associated coding analysis, as shown in Figure 6, we propose future research aiming to achieve the long-term alignment goal by identifying three important challenges from near-term to long-term objectives, including the Specification Game (Section 6.1), Dynamic Co-evolution of Alignment (Section 6.2), and Safeguarding Coadaptation (Section 6.3).
-6.1. Specification Game
+In this section, we consolidate key findings from our systematic analysis of the framework and the reviewed literature (in Figure 2.
+To accurately capture the distribution of existing research and identify current gaps and challenges, we quantified the number of relevant studies corresponding to each dimension in the Bidirectional Human-AI Alignment.
+We elaborate on key findings below.
+Underexplored Dimensions in Aligning AI with Humans.
+Current AI alignment research has primarily focused on incorporating explicitly stated human values, often gathered through direct feedback mechanisms such as ratings, rankings, or instructions.
+However, several critical dimensions remain underexplored.
+First, the use of implicit human feedback — such as behavioral cues, physiological signals, or interaction patterns—and simulated human value feedback has received limited attention, despite their potential to provide rich, context-sensitive information about human values.
+Second, most alignment efforts concentrate on model training stages, whereas developing and customizing AI models during inference or through interactive processes for embedding and evaluation values remains significantly underexplored.
+Enabling real-time adaptation of AI behavior to human input could enhance alignment in dynamic or personalized contexts.
+Third, human-in-the-loop evaluation, which involves assessing AI systems through active human participation and feedback, is rarely used in comparison to fully automated evaluation metrics.
+Expanding research into these areas is essential for advancing more robust, responsive, and context-aware alignment.
+Underexplored Dimensions in Aligning Humans with AI.
+Human-centered alignment research has predominantly emphasized designing AI systems that facilitate human understanding through sensemaking and explanation—primarily by clarifying the justifications behind AI decisions to foster user trust and reliance.
+However, this focus often overlooks the broader goal of fostering AI literacy—the essential skills and competencies individuals need to understand, critique, use, and interact effectively with AI systems.
+Despite its foundational role in responsible AI engagement, AI literacy remains an underexplored area.
+Furthermore, while numerous studies have proposed interactive mechanisms and prototypes to support human-AI collaboration, they commonly assume that AI operates in a subordinate or assistive role.
+As AI systems grow increasingly capable, research must also consider collaborative dynamics between humans and AI with equal or superior capabilities.
+Additionally, the ethical auditing of AI from a human-centered perspective and the societal-level impacts of AI — such as changes in human behavior, social relationships, and public responses—have not been sufficiently examined.
+These dimensions are critical for ensuring meaningful and equitable alignment between humans and evolving AI technologies.
+5 Near to Long-term Risks and Opportunities
+Drawing upon insights gained from the development of our framework and the associated systematic review analysis, we propose future research aiming to achieve the long-term alignment goal by identifying three important challenges from near-term to long-term objectives, including the Specification Game, Dynamic Co-evolution of Alignment, and Safeguarding Coadaptation.
+5.1 Specification Game
 An important near-term challenge is resolving the “Specification Game”, which involves precisely defining and implementing AI goals and behaviors to align with human intentions and values.
 Next, we will introduce how synergistic efforts from two directions can potentially address this challenge.
-[…]
-[…]
-Individuals often possess value systems that encompass multiple values with varying priorities, rather than a single value, to guide their behaviors (Schwartz, 1994, 2012).
+Integrate fully specified human values into aligning AI.
+Individuals often possess value systems that encompass multiple values with varying priorities, rather than a single value, to guide their behaviors schwartz1994there ; schwartz2012overview .
 Also, these priorities can change dynamically throughout an individual’s life stages.
-As such, It is more realistic to select values compatible with specific societies or situations, given the fact that we live in a diverse world (Gabriel, 2020).
-Future research, inspired by Social Choice Theory (Arrow, 2012), could focus on using democratic processes to aggregate individual values into collective agreements.
-Building on the summaries in Sections  4.1 and  5.2, researchers can employ democratic methods to identify diverse subsets of human values for AI alignment.
+As such, It is more realistic to select values compatible with specific societies or situations, given the fact that we live in a diverse world gabriel2020artificial .
+Future research, inspired by Social Choice Theory arrow2012social , could focus on using democratic processes to aggregate individual values into collective agreements.
+Building on the summaries in Sections  3.1, researchers can employ democratic methods to identify diverse subsets of human values for AI alignment.
 Additionally, creating datasets that represent these values is crucial.
 Besides, it is crucial yet challenging for AI designers to investigate how to fully specify the appropriate values and to further integrate these values into AI alignment.
-Future important area involves developing algorithms, such as the Bradley-Terry Model (Rafailov et al., 2023) or Elo Rating System (Bai et al., 2022b), to convert heterogeneous human values into AI-compatible formats for training reward models and guiding reinforcement learning.
+Future important area involves developing algorithms, such as the Bradley-Terry Model rafailov2024direct or Elo Rating System bai2022constitutional , to convert heterogeneous human values into AI-compatible formats for training reward models and guiding reinforcement learning.
 Researchers should also explore AI models capable of aligning with unstructured human data, including free-form descriptions of values, multimedia, or sensor recordings depicting human behavior.
-[…]
+Elicit nuanced and contextual human values during diverse interactions.
 Current alignment methods use instructions, ratings, and rankings to infer human values, which can not fully capture all relevant human values and constraints.
@@
 Additionally, systems that track interactions to hypothesize and validate implicit human values in real-time should be designed.
-6.2. Dynamic Co-evolution of Alignment
+5.2 Dynamic Co-evolution of Alignment
 The challenge ahead lies in comprehending and effectively navigating the dynamic interplay among human values, societal evolution, and the progression of AI technologies.
 Future studies in these directions aim to bolster a synergistic co-evolution between AI and human societies, adapting both to each other’s changes and advancements.
-[…]
-(i) Existing literature often treats AI alignment as static, ignoring its dynamic nature.
+Co-evolve AI with changes in humans and society.
+Existing literature often treats AI alignment as static, ignoring its dynamic nature.
 A long-term perspective must consider the co-evolution of AI, humans, and society.
@@
 This challenge could potentially be addressed by forecasting the potential evolution trajectories of human values or behavioral patterns, and preparing AI with the flexibility to adapt in advance, for example, through prompting or intervention strategies.
-[…]
-(i) While current AI systems lag behind humans in many tasks, identifying and handling AI mistakes, including knowing when to seek human intervention, remains essential.
+Adapt humans and society to the latest AI advancements.
+While current AI systems lag behind humans in many tasks, identifying and handling AI mistakes, including knowing when to seek human intervention, remains essential.
 Future research should focus on developing validation mechanisms that enable humans to interpret and verify AI outputs.
-This could involve designing interfaces that allow humans to request step-by-step justifications from AI or integrating tools to verify the truthfulness of AI referring to Section 4.4.1.
+This could involve designing interfaces that allow humans to request step-by-step justifications from AI or integrating tools to verify the truthfulness of AI referring to Section 3.2.
 Additionally, developing interfaces that enable groups of humans to collaboratively validate AI outputs and creating scalable validation tools for large-scale applications are important directions.
@@
 Understanding these dynamic changes is essential for grasping the broader impact of AI on humanity and society.
-6.3. Safeguarding Co-adaptation
+5.3 Safeguarding Co-adaptation
 As AI gains autonomy and capability, the risks associated with its instrumental actions, as a means toward accomplishing its final goals, increase.
@@
 We next explore future research to address this challenge from both directions.
-[…]
-(i) As advanced AI systems become more complex, they present greater challenges for human interpretation and control.
+Specify the goals of an AI system into interpretable and controllable instrumental actions for humans.
+As advanced AI systems become more complex, they present greater challenges for human interpretation and control.
 It is crucial to empower humans to detect and interpret AI misconduct and enable human intervention to prevent power-seeking AI behavior.
@@
 These components should be human-interpretable, enabling scenario testing.
-(ii) Furthermore, advanced AI systems may intentionally mislead or disobey humans, generating plausible fabrications (Johnson and Iziev, 2022).
+(ii) Furthermore, advanced AI systems may intentionally mislead or disobey humans, generating plausible fabrications johnson2022ai .
 Developing reliable interpretability mechanisms to validate the faithfulness and honesty of AI behaviors is essential.
-This includes correlating AI behaviors with internal neuron activity signals, akin to physiological indicators in human polygraph tests (Association et al., 2004).
+This includes correlating AI behaviors with internal neuron activity signals, akin to physiological indicators in human polygraph tests american2004truth .
 Inspecting these indicators can help humans assess the truthfulness of AI interpretations and prevent risky actions.
-[…]
-(i) Preventing advanced AI from engaging in risky actions requires robust human supervision.
+Empower humans to identify and intervene in AI instrumental and final strategies in collaboration.
+Preventing advanced AI from engaging in risky actions requires robust human supervision.
 Essential steps include developing training and simulation environments with scenario-based exercises and timely feedback, and creating interactive dashboards for real-time monitoring.
@@
 These tools should learn normal AI behavior and flag deviations immediately.
-Integrating training environments, interactive dashboards, and scalable diagnostic tools will enhance human ability to manage AI risks, ensuring better alignment with human values.
-7. Limitations and Conclusion
+Integrating training environments, interactive dashboards, and scalable diagnostic tools will enhance human ability to ensure better alignment with human values.
+6 Limitations
 One limitation of this work is the scope of the sampled and filtered papers.
@@
 Our surveyed papers and team members primarily focus on computing-related fields like ML, NLP, and HCI, though alignment research also involves disciplines like cognitive science, psychology, and STS (Science, Technology, and Society).
-Our framework can naturally extend to these areas as needed.
+Our framework can naturally extend to these areas.
 Despite these limitations, we believe our bidirectional human-AI alignment framework serves as a foundational reference for future researchers.
-In conclusion, this study clarifies the definitions and scope of core terminologies of human-AI alignment and conducts a systematic review of over 400 related papers spanning diverse domains such as NLP, AI, HCI, and social science.
-Additionally, we introduce a novel conceptual framework of “Bidirectional Human-AI Alignment”, structuring the surveyed literature taxonomies into “aligning AI to humans” and “aligning humans to AI” with detailed categories and example papers.
-Furthermore, we identify limitations and risks in this area quantitatively and qualitatively, analyzing a fine-grained human value taxonomy, interaction modes for alignment, and discrepancies between AI and human evaluation.
-To pave the way for future studies, we discuss five stages to achieve the alignment goals from near-term to long-term perspectives and identify new possibilities to highlight future directions and opportunities in research.
+7 Conclusion
+This study clarifies the conceptual foundations of human-AI alignment by analyzing how key terms are defined and operationalized across over 400 papers from NLP, HCI, ML, and related domains.
+We introduce the Bidirectional Human-AI Alignment framework, which organizes alignment efforts into two interdependent directions: aligning AI with humans values, and aligning humans with AI – enabling humans to effectively understand, evaluate, and adapt to AI systems.
+Our analysis identifies critical gaps in current literature, including limited support for long-term interaction, underdeveloped models of human values, and challenges in mutual intelligibility.
+We conclude with three central challenges—specification gaming, scalable oversight, and dynamic alignment—and offer actionable recommendations to support future research aimed at fostering reciprocal, robust, and context-aware approaches to human-AI alignment.
 APPENDIX
-Appendix A Systematic Literature Review
-A.1. Systematic Literature Review Process
-To understand the research literature relevant to the ongoing, mutual process of human-AI alignment, we performed a systematic literature review based on the PRISMA guideline (Page et al., 2021; Stefanidi et al., 2023).
+8 Systematic Literature Review
+8.1 Systematic Literature Review Process
+To understand the research literature relevant to the ongoing, mutual process of human-AI alignment, we performed a systematic literature review based on the PRISMA guideline Pagen71 ; 10.1145/3544548.3581332 .
 Figure 3 shows the workflow of our process for paper coding and developing the bidirectional human-AI alignment framework.
 We introduce the step details below.
-A.1.1. Identification and Screening with Keywords.
-We started with papers published in the AI-related domain venues (including NLP, HCI, and ML fields) beginning from the advent of general-purpose generative AI to present, i.e., primarily between January, 2019 and January, 2024 (see details in Appendix A.2).
+Figure 3: The selection and refinement process of our systematic literature review.
+We referred to the PRISMA guideline Pagen71 ; 10.1145/3544548.3581332 to report the workflow.
+From the identification of 34,213 records by keyword search, to screen eligible papers against our criteria and arriveg at our final corpus of 411 papers.
+For each of the stages where literature reviews were excluded (identification, screening, and eligibility) we further present the total of excluded records.
+8.1.1 Identification and Screening with Keywords.
+We started with papers published in the AI-related domain venues (including NLP, HCI, and ML fields) beginning from the advent of general-purpose generative AI to present, i.e., primarily between January, 2019 and January, 2024 (see details in Appendix 8.2).
 We retrieved 34,213 papers in the initial Identification stage.
-Further, we collectively defined a list of keywords (see details in Appendix A.3) and screened for papers that included at least one of these keywords (e.g., human, alignment) or their variations either in the title or abstract.
-We included 2,136 papers in this Screening stage.
-A.1.2. Assessing Eligibility with Criteria
+Further, we collectively defined a list of keywords (see details in Appendix 8.3) and screened for papers that included at least one of these keywords (e.g., human, alignment) or their variations in the title or abstract.
+We included 2,136 papers in Screening stage.
+8.1.2 Assessing Eligibility with Criteria
 We further filtered the 2,136 papers based on explicit inclusion and exclusion criteria, i.e., the Eligibility stage.
@@
 We included papers that could potentially answer any of these questions.
-Further, based on the scope in Section 3.4, we excluded papers that did not meet our inclusion criteria.
-This resulted in a final corpus of 411 papers, which were analyzed in detail using qualitative coding (see Appendix A.4 for more details).
-A.1.3. Qualitative Code Development.
-Referring to the code development process in Lee et al.
-(2024), we first conducted qualitative coding for each paper by identifying relevant sentences that could answer the above research questions, and entering short codes to describe them into a codebook.
+Further, based on the scope in Section LABEL:sec:scope, we excluded papers that did not meet our inclusion criteria.
+This resulted in a final corpus of 411 papers, which were analyzed in detail using qualitative coding (see Appendix 8.4 for more details).
+8.1.3 Qualitative Code Development.
+Referring to the code development process in (lee2024design, ), we first conducted qualitative coding for each paper by identifying relevant sentences that could answer the above research questions, and entering short codes to describe them into a codebook.
 We iteratively coded relevant sentences from each paper through a mix of inductive and deductive approaches, which allowed flexibility to expand, modify or change the driving research questions based on our learnings as we went through the process.
@@
 All team members coded each of their assigned papers to answer all six questions (if applicable) introduced above.
-A.1.4. Framework Development and Rigorous Coding.
+8.1.4 Framework Development and Rigorous Coding.
 After developing annotations, all authors collaborated to create the bidirectional human-AI alignment framework by integrating the annotations within each of the codes.
@@
 Additionally, we strengthened the framework by reviewing papers from the AI Ethics conferences (including FAccT and AIES), and related work of the collected papers that covered other domains such as psychology and social science.
-We further added missing codes and papers to ensure comprehensive coverage (see Appendix A.2 for details).
+We further added missing codes and papers to ensure comprehensive coverage (see Appendix 8.2 for details).
 The final bidirectional human-AI alignment framework, with detailed topologies, is presented in Section 3.
@@
 [^2] The joint probability of agreement for the paper annotations was 0.78.
-These codes were then used to perform quantitative and qualitative analyses, as presented in Section 5.
-Figure 7.
-The selection and refinement process of our systematic literature review.
-We referred to the PRISMA guideline (Page et al., 2021; Stefanidi et al., 2023) to report the workflow.
-From the identification of 34,213 records by keyword search, to screen eligible papers against our criteria and arriveg at our final corpus of 411 papers.
-For each of the stages where literature reviews were excluded (identification, screening, and eligibility) we further present the total of excluded records.
-[…]
-A.2. Venues
+These codes were then used to perform quantitative and qualitative analyses, as presented in Section LABEL:sec:analysis.
+8.2 Venues
 We primarily focused on papers from the fields of HCI, NLP, and ML ranging from year 2019 to 2024 January.
@@
   • Others: ArXiv, FAccT, AIES, and other related work
-Additionally, we also consolidate the framework by reviewing the papers published in FAccT and AIES (i.e., important venues for AI Ethics research) between 2019 and 2024 and supplemented the codes, including the AI Regulatory and Policy code in Section 4.4.2 and the exemplary paper of Regulating ChatGPT (Hacker et al., 2023)), which were not covered by the original collections.
+Additionally, we also consolidate the framework by reviewing the papers published in FAccT and AIES (i.e., important venues for AI Ethics research) between 2019 and 2024 and supplemented the codes, including the AI Regulatory and Policy code in Section LABEL:sec:impacts and the exemplary paper of Regulating ChatGPT hacker2023regulating ), which were not covered by the original collections.
 Also, we include a number of papers in the “Other” class are found by related work that are highly relevant to this topic.
-A.3. Keywords
+8.3 Keywords
 We decided on a list of keywords relevant to bidirectional human-AI alignment. The detailed keywords include:
@@
   • Ethics: Bias, Fairness
-A.4. Inclusion and Exclusion Criteria
+8.4 Inclusion and Exclusion Criteria
 To further filter the most relevant papers among the keyword-filtered 2136 papers, we identified the six most important research questions we are interested in.
@@
 Inclusion Criteria:
-  • [[[[Human values]]]] we include papers that study human value definition, specification and evaluation in AI systems.
-  • [[[[AI development techniques]]]] We include techniques of developing AI that aim to be more consistent with human values with interactions along all AI development stages (e.g., data collection, model construction, etc.)
-  • [[[[AI evaluation, explanation and utilization]]]] we include papers that build human-AI interactive systems or conduct human studies to better evaluate, explain, and utilize AI systems.
-  • [[[[building dataset with human interaction]]]] especially responsible dataset.
+  • [[Human values]] we include papers that study human value definition, specification and evaluation in AI systems.
+  • [[AI development techniques]] We include techniques of developing AI that aim to be more consistent with human values with interactions along all AI development stages (e.g., data collection, model construction, etc.)
+  • [[AI evaluation, explanation and utilization]] we include papers that build human-AI interactive systems or conduct human studies to better evaluate, explain, and utilize AI systems.
+  • [[building dataset with human interaction]] especially responsible dataset.
 Exclusion Criteria:
-  • [[[[Alignment not between human & AI]]]] we do not include alignment studies that are not between human and AI, such as entity alignment, cross-lingual alignment, cross-domain alignment, multi-modal alignment, token-environment alignment, etc.
-  • [[[[AI models beyond LLMs - Modality]]]] we do not focus on AI models other than LLMs (e.g., 3D models, VR/AR, voice assistant, spoken assistant), our primary model modality is text.
+  • [[Alignment not between human & AI]] we do not include alignment studies that are not between human and AI, such as entity alignment, cross-lingual alignment, cross-domain alignment, multi-modal alignment, token-environment alignment, etc.
+  • [[AI models beyond LLMs - Modality]] we do not focus on AI models other than LLMs (e.g., 3D models, VR/AR, voice assistant, spoken assistant), our primary model modality is text.
 Specifically, we do not consider audio / video data; we do not consider pure computer vision modality.
-  • [[[[No human-AI interaction]]]] we do not consider studies that do not involve the interaction between human and AI, such as (multi-agent) reinforcement learning.
+  • [[No human-AI interaction]] we do not consider studies that do not involve the interaction between human and AI, such as (multi-agent) reinforcement learning.
 Specifically, we do not consider interactions via voices/speech, Do not consider game interaction; Do not consider interaction for Accessibility; Do not consider Mobile interaction; Not consider autonomous vehicle interaction wearable devices, or Physical interaction;
-  • [[[[Tasks]]]] art and design, emotion.
-  • [[[[No human included]]]]
-  • [[[[focus on English]]]] primarily focus on English as the main language;
-  • [[[[Application]]]] not include the NLP papers tailored for a specific traditional task, such as translation, entity recognition, sentiment analysis, knowledge graph, adversarial and defense, topic modeling, detecting AI generations, distillation, low resource, physical robots, text classification, games, image-based tasks, hate speech detection, Human Trafficking, etc.
-  • [[[[Visualizing Embeddings]]]] Visualizing/interacting transformer embeddings?
-  • [[[[Embedding-based]]]] explanation, evaluation, etc.
-  • [[[[multi-agent reinforcement learning with self-play and population play]]]] techniques, such as self-play (SP) or population play (PP), produce agents that overfit to their training partners and do not generalize well to humans.
+  • [[Tasks]] art and design, emotion.
+  • [[No human included]]
+  • [[focus on English]] primarily focus on English as the main language;
+  • [[Application]] not include the NLP papers tailored for a specific traditional task, such as translation, entity recognition, sentiment analysis, knowledge graph, adversarial and defense, topic modeling, detecting AI generations, distillation, low resource, physical robots, text classification, games, image-based tasks, hate speech detection, Human Trafficking, etc.
+  • [[Visualizing Embeddings]] Visualizing/interacting transformer embeddings?
+  • [[Embedding-based]] explanation, evaluation, etc.
+  • [[multi-agent reinforcement learning with self-play and population play]] techniques, such as self-play (SP) or population play (PP), produce agents that overfit to their training partners and do not generalize well to humans.
 We acknowledge the extensive scope and rapid advancements of research in this area, and posit that our study offers insights that can be generalized to various modalities.
@@
 Instead, we adopt a human-centered perspective to review more than 400 key studies in this domain, focusing on delineating the framework landscape, identifying limitations, future directions, and a roadmap to pave the way for future research.
-Appendix B Author contributions
+9 Selected Paper List
+10 Alignment Goals and Human Values
+Table 2: The Goals of Alignment.
+We present the six prevailing alignment goals, associating with their Definitions (middle column), Limitations and Risks (right column).
+We consider Human Values as the main goal of alignment in this work referring to an extensive analysis and arguments in existing studies gabriel2020artificial ; stuart2014whitepaper
+10.1 A Comprehensive Taxonomy of Human Values
+This conventional theory was developed without the context of human-AI interaction, which might overlook values that need to be considered for human-AI alignment.
+Therefore, we used a bottom-up approach to extract all values studied in our collected alignment literature, mapped them onto the Schwartz Theory of Basic Values, and supplemented the theory with AI-related structure and content.
+As a result, we identified the structural relationships among human values and mapped existing literature to a fine-grained taxonomy (see Table 3).
+We supplemented the traditional theory’s four high-order value types (i.e., “Self-Enhancement”, “Openness to Change”, “Conservation”, “Self-Transcendence”) with a novel high-order value type, named “Desired Values for AI Tools” that encompasses two motivational value types (i.e., “Usability” and “Human-Likeness”).
+We further organize the relationship among these value types along two dimensions schwartz2012overview : different resources (i.e., individuals, society and interaction) and different self-intentions (i.e., self-protection against threat and self-expansion and growth).
+Furthermore, we elaborate the definitions of the 12 motivational value types and their exemplary values by mapping them to relevant human-AI alignment papers from our corpus in Table 3.
+During the process of mapping, we found: 1) value terms in empirical papers were often named differently (e.g., capability and competence), or check opposites (e.g., fairness and bias); 2) there are many values not studied in our corpus, i.e., indicated as ([Uncaptioned image]) in the Figure.
+10.2 Insights into Human Values for Alignment
+Our analysis, based on the adaptation of Schwartz’s Theory of Basic Values and our comprehensive literature review, identifies three critical findings for future research:
+Table 3: The value relations and taxonomy. We consider 5 high-order value types encompassing 12 motivational value types, indicated by their sources (e.g., individuals, society and interaction).
+Value Prioritization in AI Systems.
+Human value systems are not merely subsets of values, but ordered systems with relative priorities schwartz1994there ; schwartz2012overview .
+For instance, schwartz2012overview presented the definition for this phenomenon: “a value is ordered by importance relative to other values to form a system of value priorities.
+The relative importance of multiple values guides action….The trade-off among relevant, competing values guides attitudes and behaviors.”.
+Current AI alignment algorithms, often based on datasets of human preferences ouyang2022training ; song2024preference ; rafailov2024direct , may inadvertently prioritize majority values, potentially neglecting those of marginalized groups fleisig2023fairprism .
+Future research should address this complex interplay of values in AI systems.
+Universal vs.
+Personalized AI Values.
+While certain values are universally expected from AI (e.g., capability, equity, responsibility), others may be undesirable in specific contexts ngo2022alignment (e.g., seeking power).
+Simultaneously, AI models should be adaptable to diverse human value systems sorensen2024roadmap .
+Research is needed to develop methods for identifying appropriate value sets for specific individuals or groups, and for customizing AI to align with user values while maintaining ethical principles.
+Disparities in Value Expectations and Evaluation.
+The fundamental differences between humans and AI necessitate distinct approaches to value evaluation.
+For instance, assessing AI honesty may require mechanistic interpretability bereska2024mechanistic , a more rigorous standard than that applied to humans.
+Future studies should explore methods for evaluating and explaining AI values and calibrating human expectations accordingly.
+11 Interaction Techniques for Specifying Human Values
+Our research reveals disparities in interaction techniques for human-AI value alignment across AI-centered (NLP/ML) and human-centered (HCI) domains.
+As depicted in Figure 4, this analysis focuses on three key areas:
+Domain-Specific Interaction Techniques.
+The interaction techniques in AI-centered (NLP/ML) and Human-centered (HCI) alignment studies are often differ brath2021surveying .
+NLP/ML studies primarily utilize numeric and natural language-based techniques.
+Also, NLP/ML research explore implicit feedback to extract human hidden feedback.
+In contrast, HCI research encompasses a broader range of graphical and multi-modal interaction signals (e.g., sketches, location information) beyond text and images.
+This disparity suggests potential gaps in extracting comprehensive human behavioral information.
+Stage-Specific Interaction Techniques.
+In NLP/ML, the learning stage predominantly employs rating and ranking interactions for alignment in dataset generation.
+However, when humans use AI in the inference stage, as demonstrated in HCI research, involves more diverse user interactions.
+This discrepancy highlights the need for alignment between model development and practical deployment.
+Divergent Data Utilization.
+NLP/ML typically uses interaction outputs as training datasets, while HCI analyzes this data to understand human behavior and feedback.
+As AI systems evolve, developing new interaction modes to capture a broader spectrum of human expression becomes crucial.
+Figure 4: The interaction techniques for specifying values in human-AI alignment.
+    Takeaways of Interaction Techniques for Alignment.
+1.
+Some common human feedback styles used in NLP/ML are not often studied in HCI.
+2.
+Diverse human interactive feedback in HCI are not fully used in AI development in NLP/ML.
+12 Challenges in Achieving Alignment
+The concept of alignment in AI research has a long history, tracing back to 1960, when AI pioneer Norbert Wiener wiener1960some described the AI alignment problem as: “If we use, to achieve our purposes, a mechanical agency with whose operation we cannot interfere effectively … we had better be quite sure that the purpose put into the machine is the purpose which we really desire.” Discussion around intelligent agents and the associated concerns relating to ethics and society have emerged since then stuurman2001software .
+Next, we discuss the well-known challenges encountered in achieving alignment.
+Challenge 1: Outer and Inner Alignment.
+In the context of “intelligent agents,” until now, AI alignment research has aimed to ensure that any AI systems that would be set free to make decisions on our behalf would act appropriately and reduce unintended consequences wooldridge1999intelligent ; stuurman2001software ; russell2016artificial .
+At the near-term stage, aligning AI involves two main challenges: carefully specifying the purpose of the system (outer alignment i.e., providing well-specified rewards ngo2022alignment ) and ensuring that the system adopts the specification robustly (inner alignment, i.e., ensuring that every action given an agent in a particular state learns desirable internally-represented goals ngo2022alignment ).
+Significant efforts have been made, for inner alignment, to align AI systems to follow alignment goals of an individual or a group (e.g., instructions, preferences, values, and/or ethical principles) ouyang2022training and to evaluate the performance of alignment santurkar2023whose .
+However, for outer alignment, AI designers are still facing difficulties in specifying the full range of desired and undesired alignment goals of humans.
+Challenge 2: Specification Gaming.
+To learn human alignment goals, AI designers typically provide an objective function, instructions, reward function, or feedback to the system, which is often unable to completely specify all important values and constraints that a human intended hemphill2020human .
+Hence, AI designers resort to easy-to-specify proxy goals such as maximizing the approval of human overseers wiki_AI_alignment , which results in “specification gaming” krakovna2020specification or “reward hacking” pan2022effects issues (i.e., AI systems can find loopholes that help them accomplish the specific objective efficiently but in unintended, possibly harmful ways).
+Additionally, the black-box nature of neural networks brings additional ethical and safety concerns for alignment because humans don’t know about the inner state and the actions AI leveraged to achieve the output.
+Consequently, AI systems might make “correct” decisions with “incorrect” reasons, which are difficult to discern.
+Society is already facing these issues, such as data privacy tucker2018privacy , algorithmic bias hajian2016algorithmic , self-driving car accidents bonnefon2016social , and more.
+As a result, these considerations necessitate considering human-AI interaction in AI alignment for specification and evaluation, ranging from addressing problems around who uses an AI system, with what goals to specify, and if the AI system perform its intended function from the user’s perspective.
+Challenge 3: Scalable Oversight.
+From a long-term perspective, when advanced AI systems become more complex and capable (e.g., AGI morris2024levels ), it becomes increasingly difficult to align them to human values through human feedback.
+Evaluating complex AI behaviors applied to increasingly challenging tasks can be slow or infeasible for humans to ensure all sub-steps are aligned with their values terry2023ai .
+Therefore, researchers have begun to investigate how to reduce the time and effort for human supervision, and how to assist human supervisors, referred to as Scalable Oversight amodei2016concrete .
+Challenge 4: Dynamic Nature.
+As AI systems become increasingly powerful, the alignment solutions must also adapt dynamically since human values and preferences change as well.
+As dautenhahn2000living posit, AI systems may be neither humane nor desirable if we do not ask questions about the long-term cognitive and social effects of social agent systems (e.g., how will agent technology affect human cognition).
+All these considerations call for a long-term and dynamic perspective to address human-AI alignment as an ongoing, mutual process with the collective efforts of cross-domain expertise.
+Challenge 5: Existential Risk.
+Further, some AI researchers claim that bengio2024managing advanced AI systems will begin to seek power over their environment (e.g., humans) once deployed in real-world settings, as such behavior may not be noticed during training.
+For example, some language models seek power in text-based social environments by gaining money, resources, or social influence pan2023rewards .
+Consequently, some hypothesize that future AI, if not properly aligned with human values, could pose an existential risk to humans dafoe2016yes .
+13 Author contributions
 This project was a team effort, built on countless contributions from everyone involved.
-To acknowledge individual authors’ contributions and enable future inquiries to be directed appropriately, we followed the ACM’s policy on authorship (of Publications, 2024) and listed contributors for each part of the paper below.
-B.1. Overall Author List and Contributions
+To acknowledge individual authors’ contributions and enable future inquiries to be directed appropriately, we followed the ACM’s policy on authorship acm_authorship_2024 and listed contributors for each part of the paper.
+13.1 Overall Author List and Contributions
 Project Lead
 The project lead initialized and organized the project, coordinated with all authors, participated in the entire manuscript.
-  • Hua Shen (University of Michigan, huashen@umich.edu): Initiated and led the overall project, prepared weekly project meetings, filtered papers, designed dimensions and codes (initial, revision), coded all papers, initiated the framework and developed human value and interaction modes analysis figures, participated in drafting all sections, paper revision and polishing.
+  • Hua Shen (NYU Shanghai, New York University, huashen@nyu.edu): Initiated and led the overall project, prepared weekly project meetings, filtered papers, designed dimensions and codes (initial, revision), coded all papers, initiated the framework and developed human value and interaction modes analysis figures, participated in drafting all sections, paper revision and polishing.
 Team Leads
 The team leads organized all team events, coordinated with leads and members, contributed to a portion of manuscript.
-  • Tiffany Knearem (Google, tknearem@google.com): Led the HCI team, prepared weekly team meetings, filtered papers, designed dimensions and codes (initial, revision), coded partial papers, ideated the framework and analysis and future work content, participated in writing (Critical Thinking and AI Impact on Human sections), paper revision and polishing.
+  • Tiffany Knearem (MBZUAI, Tknearem@gmail.com): Led the HCI team, prepared weekly team meetings, filtered papers, designed dimensions and codes (initial, revision), coded partial papers, ideated the framework and analysis and future work content, participated in writing (Critical Thinking and AI Impact on Human sections), paper revision and polishing.
   • Reshmi Ghosh (Microsoft, reshmighosh@microsoft.com): Led the NLP/AI team, prepared weekly team meetings, filtered papers, coded partial papers, ideated the framework and analysis and future work content, participated in writing (AI evaluation section), paper revision and polishing.
@@
   • Kenan Alkiek (University of Michigan, kalkiek@umich.edu): filtered papers, coded partial papers, data processing and analysis, ideated paper analysis and future work, paper revision and polishing, mainly involved in NLP Team
-  • Kundan Krishna (Carnegie Mellon University, kundank@andrew.cmu.edu): filtered papers, coded partial papers, ideated the framework and future work, participated in writing (Customizing AI section), designed dimensions and codes (initial, revision), paper revision and polishing, mainly involved in NLP Team
+  • Kundan Krishna (Apple, kundank@andrew.cmu.edu): filtered papers, coded partial papers, ideated the framework and future work, participated in writing (Customizing AI section), designed dimensions and codes (initial, revision), paper revision and polishing, mainly involved in NLP Team
   • Yachuan Liu (University of Michigan, yachuan@umich.edu): filtered papers, coded partial papers, participated in writing (revised Integrate General Value and Customization content sections), paper revision and polishing, mainly involved in NLP Team
@@
   • David Jurgens (University of Michigan, jurgens@umich.edu): provided advice throughout the project, including iterative discussions on project milestones and content ideation, organized several meetings to receive feedback from external audiences, and participated in paper revision and proofreading.
-Acknowledgements.
+ACKNOWLEDGMENTS
 We thank Eric Gilbert for his constructive feedback on human-centered insights on alignment, thank Eytan Adar for his valuable guidance on designing the interaction techniques for human-AI alignment, and thank Elizabeth F.